### v4

# Batch Predictive Inference

Abstract
Constructing prediction sets with coverage guarantees for unobserved outcomes is a core problem in modern statistics. Methods for predictive inference have been developed for a wide range of settings, but usually only consider test data points one at a time. Here we study the problem of distribution-free predictive inference for a batch of multiple test points, aiming to construct prediction sets for functions—such as the mean or median—of any number of unobserved test datapoints. This setting includes constructing simultaneous prediction sets with a high probability of coverage, and selecting datapoints satisfying a specified condition while controlling the number of false claims.
For the general task of predictive inference on a function of a batch of test points, we introduce a methodology called batch predictive inference (batch PI), and provide a distribution-free coverage guarantee under exchangeability of the calibration and test data. Batch PI requires the quantiles of a rank ordering function defined on certain subsets of ranks. While computing these quantiles is NP-hard in general, we show that it can be done efficiently in many cases of interest, most notably for batch score functions with a compositional structure—which includes examples of interest such as the mean—via a dynamic programming algorithm that we develop. Batch PI has advantages over naive approaches (such as partitioning the calibration data or directly extending conformal prediction) in many settings, as it can deliver informative prediction sets even using small calibration sample sizes. We illustrate that our procedures provide informative inference across the use cases mentioned above, through experiments on both simulated data and a drug-target interaction dataset.
1 Introduction
Consider a supervised learning setting where we have a dataset $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ drawn from $P_{X,Y}=P_{X}\times P_{Y\mid X}$ on $\mathcal{X}\times\mathcal{Y}$ and a batch of new test inputs $X_{n+1},\ldots,X_{n+m}$ from $P_{X}$. Our task is to predict and make inference for the unobserved outcomes $Y_{n+1},\ldots,Y_{n+m}$, or more generally learn about the conditional distributions $P_{Y_{n+1}\mid X_{n+1}},\ldots,P_{Y_{n+m}\mid X_{n+m}}$. This setting includes both regression and classification. Beyond point predictions, it is of significant interest to construct prediction sets for various functions of the unobserved outcomes $Y_{n+1},\ldots,Y_{n+m}$. For example, given a regression function $\hat{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$ trained using a subset of the data, one might aim to construct a prediction set of the form $\widehat{C}_{n}(X_{n+1})=[\hat{\mu}(X_{n+1})-A,\hat{\mu}(X_{n+1})+A]$ for some $A>0$, which satisfies the marginal coverage guarantee $\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\}\geqslant 1-\alpha$, for a predetermined level $\alpha\in(0,1)$.
Distribution-free inference aims to achieve such inferential targets without imposing distributional assumptions on the sampling distribution $P_{X,Y}$, and dates back at least to the pioneering works of Wilks (1941), Wald (1943), Scheffe and Tukey (1945), and Tukey (1947, 1948). For example, conformal prediction (e.g., Saunders et al., 1999; Vovk et al., 1999, 2005, etc.) provides a general framework for achieving marginal coverage under exchangeability. Many recent works have explored the possibility of improving or generalizing this framework to achieve stronger targets, reduce computational costs, or enable inference with non-exchangeable data, etc, see Section 1.3. However, method development for joint inference on functions of multiple test points has been limited.
In this work, we develop methodology for distribution-free joint inference on multiple test points. This is a general problem, and includes as special case several examples of interest:
  1. 1.
    Prediction sets for multiple unobserved outcomes. Consider constructing an algorithm $\widehat{C}_{n}$ that likely obtains at least $1-\delta$ empirical coverage over the test set, i.e., $\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\delta}\right\}\geqslant 1-\alpha$. Compared to applying conformal prediction separately to individual test points to obtain $\mathbb{P}\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\}\geqslant 1-\alpha^{\prime}$, $j=1,\ldots,m$, for some $\alpha^{\prime}$, the above coverage guarantee directly states that most prediction sets cover the true outcome.
  2. 2.
    Inference on the mean. Consider predicting the mean of the test outcomes via a prediction set $\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})$ such that $\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}Y_{n+j}\in\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})}\right\}\geqslant 1-\alpha$. This problem had a range of use cases and we illustrate it in a problem of inference on the mean of counterfactual variables. When the size $m$ of the test data set is small, methods based on concentration inequalities can generally be conservative, producing wide intervals, in contrast, as we demonstrate empirically, our methods can still be informative.
  3. 3.
    Selection of of datapoints with error control. Consider selecting test datapoints in the test set whose responses satisfy a specific condition, such as $Y_{n+j}>c$ for a predetermined threshold $c$. As $(Y_{n+j})_{1\leqslant j\leqslant m}$ are unobserved, a potential approach is to construct a selection criterion based on the training and calibration data, e.g., of the form $\hat{\mu}(X_{n+j})>\hat{T}$. One possible inferential target is the control of the probability of making more than $k$ errors at level $\alpha$, i.e., $\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})>\hat{T},Y_{n+j}\leqslant c}\right\}>k}\right\}<\alpha$, where $k$ is a predetermined target error bound. This is analogous to the notion of family-wise error rate (FWER) control in multiple hypothesis testing.
We provide more details on the above examples in Section 3. The examples turn out to be special cases of the following general problem: Given the calibration data $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{1\leqslant i\leqslant n}$ and a function $g$ that takes the (multi)set of test observations as the input, construct a prediction set $\widehat{C}(\mathcal{D}_{n})$ that satisfies
$\mathbb{P}\left\{{g\left(\{(X_{n+1},Y_{n+1}),\ldots,(X_{n+m},Y_{n+m})\}\right)\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha.$
For instance, the high-probability coverage property for multiple unobserved outcomes can be achieved by taking $g$ to be a specific quantile of the non-conformity scores of the test data. More generally, we propose a batch predictive inference methodology applicable to a wide range of target functions $g$. We then explain use cases, including those described above.
Notation. We write $\mathbb{R}$ to denote the set of real numbers and $\mathbb{R}_{\geqslant 0}$ to denote the set of nonnegative reals. The set of positive integers is denoted by $\mathbb{N}$. For a positive integer $n\in\mathbb{N}$, we write $[n]$ to denote the set $\{1,2,\ldots,n\}$ and for any $a,b\in[n]$ with $a\leqslant b$ write $X_{a:b}$ to denote the vector $(X_{a},X_{2},\ldots,X_{b})^{\top}$. We will denote the all ones vector of size $m$ as $1_{m}$. For a function $f:A\rightarrow B$, We write $\mathrm{Im}(f)$ to denote the image of a function $f$, and $f\big{|}_{C}$ to denote the restriction of $f$ to $C\subset A$. For a real number $x$, we write $\lfloor x\rfloor$, $\lceil x\rceil$, and $\text{round}(x)$ to denote the floor, ceiling, and rounding of $x$ (with 1/2 rounding up) to the nearest integer, respectively. We let $a_{+}=\max\{a,0\}$ for a real number $a\in\mathbb{R}$. We denote the number of ways to choose $r$ items with replacement from $n$ items as ${}_{n}\mathrm{H}_{r}$.
We write $\sum_{i=1}^{k}p_{i}\delta_{v_{i}}$ to denote the discrete distribution with support $\{v_{1},v_{2},\ldots,v_{k}\}$ and the probability masses $(p_{1},p_{2},\ldots,p_{k})$. For a distribution $P$, we define $Q_{\tau}(P)=\inf\{t\in\mathbb{R}:\mathbb{P}_{{X\sim Q}}\left\{{X\leqslant t}\right\}\geqslant\tau\}$ and $Q_{\tau}^{\prime}(P)=\sup\{t\in\mathbb{R}:\mathbb{P}_{{X\sim P}}\left\{{X\leqslant t}\right\}\leqslant\tau\}$. For an event $E$, we write ${\mathbbm{1}}\left\{{E}\right\}$ to denote its corresponding indicator variable. All objects (sets and functions) considered will be measurable with respect to appropriate sigma-algebras (typically the Borel sigma-algebra generated by open sets), which will not be mentioned further. For a set $D$, $\mathcal{P}(D)$ denotes its power set; or the Borel sigma algebra on $D$ if that is well-defined. We write $\mathcal{N}(\mu,\sigma^{2};[a,b])$ to denote the truncated normal distribution with mean $\mu$, variance $\sigma^{2}$, and truncation set $[a,b]$.
1.1 Main contributions
Our contributions can be summarized as below:
  1. 1.
    Batch predictive inference (batch PI): We develop the batch predictive inference (batch PI): methodology for distribution-free inference on a function of multiple unobserved test outcomes. Our targets include a broad range of functions satisfying a certain monotonicity property, such as the mean or quantiles. Furthermore, we extend this approach to achieve simultaneous inference on multiple quantiles of test scores. Batch PI can provide useful inference when the calibration dataset size is comparable to—or even smaller than—the test size, a scenario in which we show that naive approaches fail.
  2. 2.
    Efficient algorithms for the batch PI procedure: We show that the batch PI procedure is generally NP-hard to compute, but it can be simplified for many target functions of practical interest, such as the mean and quantiles. For quantiles, and more generally for “sparse” functions depending only on a few quantiles, we establish how the computational burden can be reduced substantially, making the approach feasible in routine applications. For the mean, and more generally for functions satisfying a certain compositional structure, we present a polynomial-time dynamic programming algorithm for batch PI.
  3. 3.
    Use cases in statistical inference problems: We develop use cases of the batch PI methodology in various statistical inferential problems: (1) constructing simultaneous prediction sets for multiple individual outcomes, (2) selecting individuals with error control, and (3) inference on counterfactual variables. The last use case relies on a more general methodology that we develop for the setting of coverage under covariate shift.
  4. 4.
    Empirical evaluation: We empirically examine the performance of batch PI-based methods in simulations and via an illustration on a drug-target interaction dataset. The empirical results support that our procedure achieves the theoretical guarantees, and provides practically useful predictive inference.
1.2 Problem setting
We observe data points $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$, with $(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}$ for $i\in[n]$, where $\mathcal{X}$ is a feature space and $\mathcal{Y}$ is an outcome space. Here and below, sets refer to multisets, and allow repetitions of elements. We view $\mathcal{D}_{n}$ as a calibration dataset, in the sense that we assume access to another independent dataset where one can build a predictor $\hat{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$. The calibration dataset $\mathcal{D}_{n}$ is used to construct a prediction set, leveraging $\hat{\mu}$. The prediction set is applied to the test dataset, which consists of $m\geqslant 1$ observed test features $X_{n+1},X_{n+2},\ldots,X_{n+m}$ without their corresponding outcomes $Y_{n+1},Y_{n+2},\ldots,Y_{n+m}$. We denote each data point as $Z_{i}=(X_{i},Y_{i})$, for $i\in[n+m]$.
Given a function $g:\{Z_{n+1},\ldots,Z_{n+m}\}\mapsto g(\{Z_{n+1},\ldots,Z_{n+m}\})$ of interest of the set of test points $\{Z_{n+1},\ldots,Z_{n+m}\}$, our goal is to construct a prediction set for the unobserved value $g(\{Z_{n+1},\ldots,Z_{n+m}\})$; which depends on the unobserved outcomes $Y_{n+1},Y_{n+2},\ldots,Y_{n+m}$. Specifically, we aim to construct a procedure $\widehat{C}:(\mathcal{X}\times\mathcal{Y})^{n}\rightarrow\mathcal{P}(\mathbb{R})$ such that
$\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$ (1)
holds for a predefined level $\alpha\in(0,1)$. We are interested in a general setting where $m$ is not necessarily significantly smaller than the calibration set size $n$ (in contrast to cases with trivial solutions, as we will describe later), and may even be larger. We will argue that this setting covers a wide range of important scenarios.
We now need some notations: For any vector $v\in\mathbb{R}^{m}$, let $v_{\uparrow}=(v_{(1)},\ldots,v_{(m)})$ be the vector $v$ sorted in a non-decreasing order, so that $v_{(1)}\leqslant\ldots\leqslant v_{(m)}$. For $z\in(\mathcal{X}\times\mathcal{Y})^{m}$ and a “score” function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, define $s(z)=(s(z_{1}),s(z_{2}),\ldots,s(z_{m}))$ by element-wise application of $s$. We denote $S_{i}=s(Z_{i})$ for all $i\in[m+n]$. Let $\mathbb{R}^{m}_{\uparrow}=\{x\in\mathbb{R}^{m}:x_{1}\leqslant x_{2}\leqslant\ldots\leqslant x_{m}\}$ be the set of monotone non-increasing vectors. For two vectors $u=(u_{1},\ldots,u_{d})^{\top},v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$, we write $u\preceq v$ if $u_{i}\leqslant v_{i}$ for all $i=1,2,\ldots,d$.
We require the following structural monotonicity condition for the target function $g$.
Condition 1 (Monotonicity of the target function).
There is a batch score function [^2] Let $\inf s=\inf\{s(x,y):(x,y)\in\mathcal{X}\times\mathcal{Y}\}$ and $\sup s=\sup\{s(x,y):(x,y)\in\mathcal{X}\times\mathcal{Y}\}$. When $s$ is unbounded, we will also need the function $h$ to be defined for all values $s_{1}\leqslant\ldots\leqslant s_{m}$ such that $s_{i}\in(\inf s,\sup s)$ for all $i\in\{2,\ldots,m-1\}$ and either $s_{1}=\inf s$ or $s_{m}=\sup s$. We will define $h(-\infty,s_{2},\ldots,s_{m})=-\infty$ if $s_{1}=\inf s=-\infty$, and $h(s_{1},\ldots,s_{m-1},\infty)=\infty$ if $s_{m}=\sup s=\infty$. $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$ and a (non-batch, per-datapoint) score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, such that
$g(\{z_{1},\ldots,z_{m}\})=h(s(z)_{\uparrow})$ (2)
holds for all $z\in(\mathcal{X}\times\mathcal{Y})^{m}$. Moreover, the function $h$ is monotone non-decreasing with respect to each coordinate, i.e.,
for $\textnormal{for any }v,\tilde{v}\in\mathbb{R}^{m}\textnormal{ with }v\preceq\tilde{v},\textnormal{ we have }h(v_{\uparrow})\leqslant h(\tilde{v}_{\uparrow}).$ (3)
Condition 1 covers a broad range of targets, from the mean $h(s_{1},\ldots,s_{m})=\frac{s_{1}+\ldots+s_{m}}{m}$ and the $q$-th quantile $h(s_{1},\ldots,s_{m})=s_{(\lceil(qm)\rceil)}$, $q\in(0,1)$, to more general targets such as the truncated mean or the proportion of scores exceeding a certain threshold. In many settings, $h$ represents a fixed function of interest, while $s$ is typically constructed using a separate dataset. For instance, in regression tasks, we can consider nonconformity scores such as $s:(x,y)\mapsto|y-\hat{\mu}(x)|$, where $\hat{\mu}$ is fitted on a separate dataset.
If the cardinality of $\mathcal{X}\times\mathcal{Y}$ is at most that of $\mathbb{R}$ (which holds for instance in the typical case that $\mathcal{X}\subset\mathbb{R}^{d}$ for some positive integer $d\geqslant 0$, and $\mathcal{Y}$ is either discrete or $\mathbb{R}$), then there is an injective map $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. Let $\mathcal{S}\subset\mathbb{R}$ be the image of $s$. Then, for any $v\in\mathbb{R}^{m}_{\uparrow}\cap\mathcal{S}^{m}$, we can define $h(v)=g(\{s^{-1}(v_{1}),\ldots,s^{-1}(v_{m})\})$. For $v\in\mathbb{R}^{m}_{\uparrow}\setminus\mathcal{S}^{m}$, we can define $h(v)$ arbitrarily. Since $s$ is injective, $h$ is well-defined. Moreover, (2) holds by definition. Thus, if the cardinality of $\mathcal{X}\times\mathcal{Y}$ is at most that of $\mathbb{R}$, (2) holds, and only the monotonicity property (3) imposes a condition.
1.3 Related work
The idea of distribution-free prediction sets dates back at least to the pioneering works of Wilks (1941), Wald (1943), Scheffe and Tukey (1945), and Tukey (1947, 1948). Distribution-free inference has been extensively studied in recent works (see, e.g., Saunders et al., 1999; Vovk et al., 1999; Papadopoulos et al., 2002; Vovk et al., 2005; Vovk, 2013a; Lei et al., 2013; Lei and Wasserman, 2014; Lei et al., 2018; Angelopoulos et al., 2023; Guan, 2023; Romano et al., 2020; Liang et al., 2023; Dobriban and Yu, 2023). Predictive inference methods (e.g., Geisser, 2017, etc) have been developed under various assumptions (see, e.g., Bates et al., 2021; Park et al., 2022a, b; Sesia et al., 2023; Qiu et al., 2023; Li et al., 2022; Kaur et al., 2022; Si et al., 2024; Lee et al., 2024). Overviews of the field are provided by Vovk et al. (2005); Shafer and Vovk (2008), and Angelopoulos et al. (2023). For exchangeable data, conformal prediction and split conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002) provide a general framework for distribution-free predictive inference.
Distribution-free predictive inference for multiple test points has been extensively studied in the context of outlier detection and selection (Bates et al., 2023; Jin and Candès, 2023b, a; Gui et al., 2024). These works apply multiple testing methods to conformal p-values for inference on multiple test outcomes. Vovk (2013c) discuss transductive conformal methods for constructing a prediction region for the vector of test outcomes. Lee et al. (2024) introduces a method for constructing simultaneous prediction sets for multiple outcomes under covariate shift with a conditional guarantee.
Gazin et al. (2024) study a closely related problem setting to our paper, transductive conformal inference with adaptive scores. In this scenario, they derive the joint distribution of multiple test conformal p-values in the case of no ties between non-conformity scores, which is equivalent to the joint distribution of their ranks, and which we use in the proof of our Theorem 1. Gazin et al. (2024) also give intriguing equivalent characterizations of this distribution, for instance in terms of Pólya urns (see also Gazin (2024) for a functional CLT for the coverage). Further, they apply these results to several problems, including controlling the false coverage rate of the prediction sets for multiple test points. For this probleem Marques F. (2025) derived the distribution of the coverage. This problem is also considered one of our use cases in this work, and we will provide further discussion in Section 3.1.
In Section 2.4, we discuss how our procedure can be applied to situations involving covariate shift. This is relevant in light of the recent literature, which has shown significant interest in extending the conformal prediction framework to handle non-exchangeable data. For instance, Tibshirani et al. (2019) proposes weighted conformal prediction for predictive inference under covariate shift, and their method is further developed in works such as Lei and Candès (2021); Candès et al. (2023), and Guan (2023). Qiu et al. (2023) and Yang et al. (2023+) introduce adaptive prediction methods with unknown covariate shift. Barber et al. (2023) introduces a robust conformal prediction approach for non-exchangeable data. Other works have explored applying the conformal prediction framework to structured datasets. For example, Dunn et al. (2022); Lee et al. (2023), and Duchi et al. (2024) provide conformal-type methods for data with a hierarchical structure, while Dobriban and Yu (2023) provides a method for data with group symmetries.
2 The batch PI methodology
Here and below, we will suppose that the calibration and test data $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1}),\ldots,(X_{n+m},Y_{n+m})$ are exchangeable, unless explicitly specified otherwise. If $m=1$, i.e., we have only one test point, then the coverage guarantee (1) can be achieved simply by standard distribution-free prediction methods, such as full and split conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002), for any function $g$. For example, if we set $g(z)$ as the nonconformity score, i.e., $g({z})=|y-\hat{\mu}(x)|$, for all $z=(x,y)$, then the condition (1) is equivalent to $\mathbb{P}\left\{{s(X_{n+1},Y_{n+1})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$, and split conformal prediction attains this guarantee with $\widehat{C}(\mathcal{D}_{n})=\Big{(}-\infty,Q_{1-\alpha}\left(\sum_{i=1}^{n}\frac{1}{n+1}\delta_{s(X_{i},Y_{i})}+\frac{1}{n+1}\delta_{\infty}\right)\Big{]}$ (Saunders et al., 1999; Vovk et al., 1999, 2005; Papadopoulos et al., 2002).
However, for multiple test points, it turns out that constructing a useful distribution-free prediction set that satisfies (1) is a nontrivial task. One can imagine a number of direct approaches (See Appendix A for details). For instance, one could partition the calibration dataset into non-overlapping subsets of the same size as the test data, and then use standard conformal prediction for the function $g$ applied to the partitions and the test data. However, this approach produces nontrivial or short prediction sets only when the calibration set size is much larger than the test set size. For example, if $n<m(1/\alpha-1)$, it leads to a trivial prediction set (See Section A.1). A direct extension of split conformal prediction has similar issues (See Section A.2). Finally, a direct extension of full conformal prediction is computationally prohibitive (See Section A.3). Thus, none of these direct approaches are practically viable in the setting we are interested in and therefore will not be given further consideration.
2.1 Proposed method: batch PI
In this section, we introduce our new batch PI procedure, which can be less conservative and more computationally efficient than the naive methods described above. To introduce our method, it is helpful to review the idea of split conformal prediction. Suppose we have only one input $X_{n+1}$ beyond the calibration data $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$. The first step is to construct a nonconformity score $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{\geqslant 0}$; based on data that is independent of the calibration and test datasets. Let us write $S_{i}=s(X_{i},Y_{i})$ for $i\in[n+1]$. The split conformal prediction set is given by
$\widehat{C}_{n}(x)=\left\{y\in\mathcal{Y}:s(x,y)\leqslant\textnormal{$\lceil(1-\alpha)(n+1)\rceil$-th smallest value of $S_{1},S_{2},\ldots,S_{n}$}\right\}.$ (4)
It is known that if $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$ are exchangeable, the prediction set $\widehat{C}_{n}$ from (4) satisfies the following coverage guarantee (Vovk et al., 2005): $\mathbb{P}\left\{{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})}\right\}\geqslant 1-\alpha$.
The key intuition is as follows: Let $S_{(1)},S_{(2)},\ldots,S_{(n)}$ be the order statistics of $S_{1},S_{2},\ldots,S_{n}$, breaking ties uniformly at random. Then, the rank $R\in\{1,\ldots,n+1\}$ such that $S_{(R)}$ is the smallest upper bound among the observed scores for the unobserved score $S_{n+1}$ follows a uniform distribution over $\{1,\ldots,n+1\}$. (where we define $S_{(n+1)}=+\infty$) Then, because $Y_{n+1}\in\widehat{C}_{n}(X_{n+1})$ is implied by $R\leqslant\lceil(1-\alpha)(n+1)\rceil$, the coverage probability is at least $1-\alpha$.
The batch PI method.
We now consider the setting of multiple test points (test size $m\geqslant 1$). Since we will need to consider not just one rank, but rather the ranks of all the test data points among the $n$ calibration data points, we define the set $H$ of monotone non-decreasing sequences of length $m$, of positive integers between one and $n+1$ as
$H$ =$=\left\{r_{1:m}:1\leqslant r_{1}\leqslant\ldots\leqslant r_{m}\leqslant n+1\right\}.$ (5)
Note that $|H|=\binom{n+m}{m}$. This set will represent the ranks of the test data points among the calibration data points.
Moreover, we also need a way to order these ranks. In the standard conformal case where $m=1$, the ranks are ordered as $11\leqslant\ldots\leqslant n+$ 1, but for our case there is no default ordering. Hence to allow for the maximum flexibility, we introduce a general rank-ordering function $\tilde{h}:H\rightarrow\mathbb{R}$ that we will use to prioritize the ranks. We will later discuss at length the choice of this function.
Given the rank-ordering function $\tilde{h}:H\rightarrow\mathbb{R}$, as well as lower and upper error levels $\beta,\gamma\in[0,1]$ satisfying $\beta+\gamma=\alpha$, we consider the following two quantiles of the distribution of the rank-ordering function given a uniform distribution over the set $H$,
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\right),\quad q_{U}=Q_{1-\gamma}\left(\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\right).$ (6)
By definition, if $R_{1:m}$ is distributed uniformly over $H$, then $\mathbb{P}\left\{{\tilde{h}(R_{1:m})\in[q_{L},q_{U}]}\right\}\geqslant 1-\alpha$. However, since we are interested in covering the values of the function $h$ (or equivalently $g$), we also need a way to define an appropriate range of $h$ values. We do this by first considering the pre-image of $[q_{L},q_{U}]$ under $\tilde{h}$, and then considering its image under $h$. It turns out that we also need to consider certain corner cases (e.g., when the rank is $n+1$), and so with $S_{(0)}=\inf s$ and $S_{(n+1)}=\sup s$, we define
$\begin{split}B_{L}&=\min\left\{h(S_{(r_{1}-1)},\ldots,S_{(r_{m}-1)}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\geqslant q_{L}\right\},\\ B_{U}&=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\leqslant q_{U}\right\}.\end{split}$ (7)
Then we construct the batch predictive inference (batch PI) prediction set as
$\widehat{C}(\mathcal{D}_{n})=\big{[}B_{L},B_{U}\big{]}.$ (8)
See Algorithm 1. For completeness, we also provide a one-sided version of the batch PI prediction set algorithm, which simplifies slightly, see Algorithm 8 in Section B.1. The validity of batch PI is proved in Theorem 1.
Input: Calibration data $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$. Score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. Test set size $m$. Batch score function $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$. Rank-ordering function $\tilde{h}:\mathbb{N}^{m}\rightarrow\mathbb{R}$. Target coverage level $1-\alpha\in[0,1]$. Lower and upper error levels $\beta,\gamma\in[0,1]$ satisfying $\beta+\gamma=\alpha$
Step 1: With $H=\left\{r_{1:m}:=(r_{1},\ldots,r_{m})^{\top}:1\leqslant r_{1}\leqslant\ldots\leqslant r_{m}\leqslant n+1\right\}$, compute the sample quantiles induced by the rank-ordering function $\tilde{h}$:
$q_{L}=Q_{\beta}^{\prime}\bigg{(}\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\bigg{)},\quad q_{U}=Q_{1-\gamma}\bigg{(}\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\bigg{)}.$
Step 2: Compute the scores $S_{i}=s(X_{i},Y_{i})$ for $i=1,2,\ldots,n$.
Step 3: Compute the bounds, with $S_{(0)}=\inf s$, and $S_{(n+1)}=\sup s$:
$B_{L}$ $=\min\left\{h(S_{(r_{1}-1)},\ldots,S_{(r_{m}-1)}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\geqslant q_{L}\right\},$
$B_{U}$ $=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\leqslant q_{U}\right\}.$
Return: Prediction set $\widehat{C}(\mathcal{D}_{n})=\big{[}B_{L},B_{U}\big{]}$
Algorithm 1 Batch Predictive Inference (batch PI)
Theorem 1 (Validity of batch PI).
Suppose that Condition 1 holds, and that the data points $Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}$ are exchangeable. Then the batch PI prediction set from (8) satisfies $\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$.
The proof is deferred to Appendix E, and here we offer some intuition. Suppose the scores $S_{1},\ldots,S_{n+m}$ are distinct almost surely, and define $R_{n+1},R_{n+2},\ldots,R_{n+m}$ as $R_{n+j}=\min\{r\in\{1,2,\ldots,n\}:S_{(r)}\geqslant S_{n+j}\},\textnormal{ for}j=1,2,\ldots,m$, where we let $R_{n+j}=n+1$ if $S_{(n)}<S_{n+j}$. Let $R_{(n+1)},\ldots,R_{(n+m)}$ be their order statistics. Through the exchangeability condition, it follows that $(R_{(n+1)},\ldots,R_{(n+m)})\sim\textnormal{Unif}(H)$ (Gazin et al., 2024). This implies that for any fixed subset $I$ of $H$ with $|I|\geqslant(1-\alpha)|H|,$ we have $\mathbb{P}\left\{{(R_{(n+1)},\ldots,R_{(n+m)})\in I}\right\}\geqslant 1-\alpha$. By construction, the set $\{r_{1:m}\in H,\tilde{h}(r_{1:m})\in[q_{L},q_{U}]\}$ satisfies this condition, and thus we have
$\mathbb{P}\left\{{h(S_{(n+1)},\ldots,S_{(n+m)})\leqslant[B_{L},B_{U}]}\right\}\geqslant\mathbb{P}\left\{{h(S_{R_{(n+1)}},\ldots,S_{R_{(n+m)}})\leqslant[B_{L},B_{U}]}\right\}\geqslant 1-\gamma,$
as desired. While the fully rigorous proof follows a similar argument, it requires more elaborate reasoning.
$\tilde{h}(1,1)$$\tilde{h}(1,2)$$\tilde{h}(2,2)$$\tilde{h}(1,3)$$\tilde{h}(1,4)$$\tilde{h}(2,3)$$\tilde{h}(2,4)$$\tilde{h}(3,3)$$\tilde{h}(3,4)$$\tilde{h}(4,4)$$h(S_{(1)},S_{(1)})$$h(S_{(1)},S_{(2)})$$h(S_{(2)},S_{(2)})$$h(S_{(1)},S_{(3)})$$h(S_{(1)},S_{(4)})$$h(S_{(2)},S_{(3)})$$h(S_{(2)},S_{(4)})$$h(S_{(3)},S_{(3)})$$h(S_{(3)},S_{(4)})$$h(S_{(4)},S_{(4)})$$B$$\infty$- ∞$\widehat{C}(\mathcal{D}_{n})$((((]]]]
Figure 1: An illustration of the batch PI method with $n=3$ calibration data points, $m=2$ test data points, and coverage $0.91-\alpha=$ 0.9. Here we show hypothetical (arbitrarily chosen) values for $\tilde{h}$ and $h$. The values of $h$ shown satisfy the monotonicity constraint from Assumption 1, which for pairs $1\leqslant i\leqslant j\leqslant 4$ and $1\leqslant k\leqslant l\leqslant 4$ reduces to $h(S_{(i)},S_{(j)})\leqslant h(S_{(k)},S_{(l)})$ if $i\leqslant j$ and $k\leqslant l$. The value $q$ is defined as the $(1-\alpha)$-th quantile of the $\tilde{h}$ values. The value $B$ is defined as the maximum of the $h$ values to the “left” of $q$. Then the batch PI prediction set is $\widehat{C}(\mathcal{D}_{n})=\big{(}-\infty,B\big{]}$, and is shown on the left.
While batch PI offers valid coverage, computing it requires finding the quantiles $q_{L},q_{U}$, as well as the interval endpoints $B_{L},B_{U}$. Specifically, the procedure includes the following computations:
  1. 1.
    $q_{L}$ and $q_{U}$ involves the computation of $\tilde{h}(r_{1},\ldots,r_{m})$ for $\binom{n+m}{m}$ elements in $H$.
  2. 2.
    $B_{L}$ and $B_{U}$ involves the computation of $h(S_{(r_{1})},\ldots,S_{(r_{m})})$ for $\lceil(1-\alpha)\binom{n+m}{m}\rceil$ rank vectors.
Since $\binom{n+m}{m}\sim(1+n/m)^{m}$, the computational cost of a direct enumeration-based approach for batch PI can be extremely high when the number of the calibration and test data points are large.
To confirm that this computation is indeed hard in general, we take the perspective of standard computational complexity theory (e.g., Garey and Johnson, 1979), where the difficulty of problems is characterized according to the number of steps it takes to execute them on a standard model of computation called the Turing machine. Tractable problems usually have a polynomial running time, while there is a potentially broader class of problems—called NP—whose solutions can be verified in polynomial time. There is a large set of difficult combinatorial problems— called NP-hard problem—that are at least as hard as any problem in NP. By showing that solving the prediction set problem can be used to solve the so-called vertex cover problem (e.g., Garey and Johnson, 1979), we show that computing batch PI is in general NP-hard.
Proposition 1 (NP-hardness of Batch PI).
Computing $B_{L}$ and $B_{U}$ in (7) is NP-hard (as a function of $n$) for general functions $h,\tilde{h}$, even when $n=m$.
However, we will show in the remainder of the paper that the computation can often be simplified at a feasible computational cost for target functions $h$ of practical interest: functions of a small number of quantiles (including single quantiles) and functions with a compositional structure.
Remark 1 (Choice of the rank ordering function).
For the choice of the rank ordering function $\tilde{h}$, we have the following considerations. To ensure validity, this function cannot depend on the calibration scores $S_{1},\ldots,S_{n}$. However, to obtain a short and informative prediction sets, the values $h(S_{(r_{1})},\ldots,S_{(r_{m})})$ when varying $(r_{1},\ldots,r_{m})$ should be similarly ordered as the values $\tilde{h}(r_{1},\ldots,r_{m})$. We will describe below two heuristic strategies to achieve this goal.
Strategy 1: Rank-ordering functionally identical to the batch score. In many settings, a simple choice would be to set $\tilde{h}=h\big{|}_{H}$, namely the restriction of the batch score function to the set of ranks (if that restriction is well defined). For instance, if we are interested in the mean of test scores, i.e., $h(s_{1},\ldots,s_{m})=\frac{1}{m}\sum_{j=1}^{m}s_{j}$, then one choice would be to set $\tilde{h}(r_{1},\ldots,r_{m})=\frac{1}{m}\sum_{j=1}^{m}r_{j}$. This ensures that the mean of the scores corresponding to a “smaller” rank vector tends to be smaller than that corresponding to a “larger” rank vector.
Strategy 2: Rank ordering based on independent split. Another approach is to use a split $\tilde{Z}_{1},\ldots,\tilde{Z}_{n}$ of the data to construct $\tilde{S}_{1}=s(\tilde{Z}_{1}),\ldots,\tilde{S}_{n}=s(\tilde{Z}_{n})$ with the same distribution as $S_{1},\ldots,S_{n}$ from the remaining split (which will be used in the batch PI procedure). Then we can consider the rank-ordering function defined as $\tilde{h}(r_{1},\ldots,r_{m})=h(\tilde{S}_{(r_{1})},\ldots,\tilde{S}_{(r_{m})})$.
2.2 Computationally tractable examples of batch PI
We now turn to discussing how the batch PI procedure simplifies to become computationally tractable in special cases of interest.
2.2.1 Inference on a quantile
Given $\delta\in(0,1)$, consider forming a prediction set for the $(1-\delta)$-th sample quantile of the unobserved scores $S_{n+1},\ldots,S_{n+m}$,
$S_{(\zeta)}^{\textnormal{test}}=\textnormal{$\zeta$-th smallest value in $(S_{n+1},S_{n+2},\ldots,S_{n+m})$, where $\zeta=\lceil(1-\delta)m\rceil$}.$
This corresponds to the batch score function $h:(s_{1},s_{2},\ldots,s_{m})\mapsto s_{\zeta}$ in Condition 1.
Observe that for this special case, we have full access to the ordering of $h$ values without knowing the exact score values, i.e., we know $S_{(i_{1})}\leqslant S_{(i_{2})}$ when $i_{1}\leqslant i_{2}$, even if the actual values of $S_{(i_{1})}$ and $S_{(i_{2})}$ are unknown. Therefore, denoting by $r_{(\zeta)}$ the $\zeta$-th smallest element in $r=(r_{1},\ldots,r_{m})$, we can set $\tilde{h}(r_{1},\cdots,r_{m})=r_{(\zeta)}$. This choice of $\tilde{h}$ recovers the exact ordering of $h$ values, in the sense that $h(S_{(r_{1})},\ldots,S_{(r_{m})})\leqslant h(S_{(r^{\prime}_{1})},\ldots,S_{(r^{\prime}_{m})})$ if and only if $\tilde{h}(r_{1},\cdots,r_{m})\leqslant\tilde{h}(r^{\prime}_{1},\cdots,r^{\prime}_{m})$. Thus, as per our discussion from Remark 1, this choice of $\tilde{h}$ is “optimal” in a sense.
Then, $q_{L}$ and $q_{U}$ in (6) can be expressed as
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\cdot\delta_{r_{(\zeta)}}\right),\qquad q_{U}=Q_{1-\gamma}\left(\sum_{r=r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\cdot\delta_{r_{(\zeta)}}\right).$
Further, since |$|\{r\in H:r_{(\zeta)}=k\}|=_{k}\mathrm{H}_{\zeta-1}\cdot_{n-k+2}\mathrm{H}_{m-\zeta}=\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}$ for $k\in[n+1]$, we have the following explicit expressions:
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k}\right),q_{U}=Q_{1-\gamma}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k}\right).$ (9)
Next, observe that $B_{U}$ in (7) for this setting can be simplified as
$B_{U}=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\tilde{h}(r_{1:m})\leqslant q_{U}\right\}=\max\left\{S_{r_{(\zeta)}}:r\in H,r_{(\zeta)}\leqslant q_{U}\right\}=S_{(q_{U})},$
and similarly $B_{L}=S_{(q_{L}-1)}$. Therefore, batch PI reduces to the following $(1-\alpha)$-prediction set for $S_{(\zeta)}^{\textnormal{test}}$:
$\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})=\left[S_{(q_{L}-1)},S_{(q_{U})}\right].$ (10)
Corollary 1 (Batch PI for quantiles).
If the data points $Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}$ are exchangeable, the prediction set $\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})$ from (9) and (10) satisfies $\mathbb{P}\left\{{S_{(\zeta)}^{\textnormal{test}}\in\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$.
For this procedure, the computational cost arises only from computing $q_{L}$ and $q_{U}$, and is relatively low since they are quantiles of discrete distributions with support size $n+1$.
Input: Calibration data $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$. Score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. Test set size $m$. Target quantile level $1-\delta\in(0,1)$. Target coverage level $1-\alpha\in[0,1]$. Lower and upper error levels $\beta,\gamma\in[0,1]$ satisfying $\beta+\gamma=\alpha$.
Step 1: With $\zeta=\lceil(1-\delta)m\rceil$, compute the sample quantiles:
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k-1}\right),\quad q_{U}=Q_{1-\gamma}\left(\sum_{k=1}^{n+1}\frac{\binom{k+\zeta-2}{\zeta-1}\binom{n+m-k-\zeta+1}{m-\zeta}}{\binom{n+m}{m}}\cdot\delta_{k}\right).$
Step 2: Compute the scores $S_{i}=s(X_{i},Y_{i})$ for $i=1,2,\ldots,n$.
Return: Prediction set $\widehat{C}^{\textnormal{bPI-q}}(\mathcal{D}_{n})=\left[S_{(q_{L})},S_{(q_{U})}\right]$
Algorithm 2 Batch PI for Inference on a Quantile
In Appendix C, we extend the above idea to describe the simplification of the batch PI procedure for general sparse functions $h$, where $h(s_{1},\cdots,s_{m})$ depends only on a small number of the $s_{j}$s.
2.2.2 Inference on the mean and general compositionally structured functions
In this section, we show how to compute the batch predictive inference prediction sets efficiently in a general setting where the rank ordering and batch score functions have a certain compositional structure, a setting that includes the important case of the mean. Recall that for a given rank-ordering function $\tilde{h}:\mathbb{N}^{m}\rightarrow\mathbb{R}$, the computation of $q_{L},q_{U}$ from (6) requires, for all $k\in\text{range}(\tilde{h})$, that we compute the number of $r_{1:m}\in H$, such that $\tilde{h}(r_{1:m})=k$.
To introduce our algorithm and ideas, let us first consider the simpler case where function $\tilde{h}$ is the sum, $\tilde{h}(r_{1:m})=\sum_{j=1}^{m}r_{j}$, for all $r_{1:m}\in H$. This is equivalent to the mean, up to scale. In that case, the problem becomes to find the number—denoted $C_{m,n,k}$—of the positive integer solutions $r_{1:m}=(r_{1},\ldots,r_{m})$ to the equation $r_{1}+r_{2}+\ldots+r_{m}=k$ with $1\leqslant r_{1}\leqslant\ldots\leqslant r_{m}\leqslant n$. Once we have $C_{m,n,k}$, we can simplify $q_{U}$ to
$q_{U}=Q_{1-\gamma}\left(\sum_{k\in\text{range}(\tilde{h})}\frac{C_{m,n,k}}{\binom{n+m}{m}}\delta_{k}\right).$ (11)
The key idea is to compute the numbers $C_{m,n,k}$ via dynamic programming. In particular, we find a recursion on the number of solutions depending on the value of $r_{m}\in[n]$. Consider any $a\in\{1,\ldots,n\}$. For a solution $r_{1:m}$, if $r_{m}=a$, then $r_{1}+\ldots+r_{m-1}=k-a$. By definition, there are $C_{m-1,n,k-a}$ such solutions. Thus, by considering all possible values of $a$ for $r_{m}\in[n]$, we obtain the recursion $C_{m,n,k}=\sum_{a=1}^{n}C_{m-1,n,k-a}.$
This can be used to develop a dynamic programming algorithm to find $C_{m,n,k}$, by sequentially computing all relevant values of $C_{1,*,*}$, $C_{2,*,*}$, …, $C_{m,*,*}$, where * denotes omitted indices. The initial conditions are: $C_{1,n,k}=I(1\leqslant k\leqslant n)$. Further, note that $C_{m,n,k}=0$ if $k<m$; because we need $m\leqslant r_{1}+r_{2}+\ldots+r_{m}=k$ to have a solution. Hence, the range or the recursion can be limited slightly, see Algorithm 3. We also point out that $\sum_{a=1}^{b}\cdot$ is considered to be zero when $b<1$.
  Input: number of summands $m$, maximum summand $n$, target sum $k$
  Initialize $C_{1,\tilde{n},\tilde{k}}=I(1\leqslant\tilde{k}\leqslant\tilde{n})$ for $\tilde{n}\in[n],\tilde{k}\in[k]$
  for $\tilde{m}=2$ to $m$ do
     for $\tilde{k}=1$ to $k$ do
        for $\tilde{n}=1$ to $n$ do
           $C_{\tilde{m},\tilde{n},\tilde{k}}\leftarrow\sum_{a=1}^{\min(\tilde{n},\tilde{k}-\tilde{m}+1)}C_{\tilde{m}-1,a,\tilde{k}-a}$
        end for
     end for
  end for
  Output: $C_{m,n,k}$
Algorithm 3 Computation of $C_{m,n,k}$ when the rank-ordering function $\tilde{h}$ is the sum
The running time of this algorithm is $O(mkn^{2})$ flops, due to a triple loop (each ranging from unity to $m,k,n$, respectively) and as the innermost computation takes $O(n)$ steps. Thus, since the range of $\tilde{h}$ ranges between $m$ and $(n+1)m$, computing $q_{U}$ from (11) by computing $C_{m,n,k}$ for all $k\in\text{range}(\tilde{h})$ has complexity $O(m^{2}kn^{3})$. [^3] Alternatively, for even faster computation with moderate sample sizes, one can estimate the quantiles $q_{L}$ and $q_{U}$ using sample quantiles. Specifically, drawing a sample from $H$ is equivalent to drawing $m$ samples from a uniform distribution over $[n+1]$ with replacement, allowing us to construct samples of $\tilde{h}(r_{1:m}),r_{1:m}\sim\text{Unif}(H)$. This approach leads to an accurate estimate of $q_{L}$ and $q_{U}$ if a sufficient number of samples is drawn.
More generally, suppose that for all $r\geqslant 1$, there is a strictly increasing function $\tilde{\Gamma}(\cdot;r):\{0,1,\ldots\}\to\{0,1,\ldots\}$ such that for any $\kappa\geqslant 1$,
$\tilde{h}(r_{1:\kappa})=\tilde{\Gamma}(\tilde{h}(r_{1:(\kappa-1)});r_{\kappa}).$ (12)
Here the function $\tilde{\Gamma}$ could be made to depend on $\kappa$, i.e., having $\tilde{h}(r_{1:\kappa})=\tilde{\Gamma}_{\kappa}(\tilde{h}(r_{1:(\kappa-1)});r_{\kappa}),$ but we omit this for simplicity. For instance, for our previous example of the sum $\tilde{h}(s_{1:\kappa})=\sum_{j\in[\kappa]}s_{j}$, we can take $\tilde{\Gamma}(a;r)=a+r,$ for all positive integers $\kappa,r,a$. Then the same reasoning by partitioning on the possible values of $r_{m}$ yields that $C_{m,n,k}=\sum_{a=1}^{n}C_{m-1,a,\tilde{\Gamma}^{-1}(k;a)}$, where $\tilde{\Gamma}^{-1}(\cdot;a)$ denotes the inverse of the function $x\mapsto\tilde{\Gamma}(x;a)$. Here, the understanding is that if the equation $\tilde{\Gamma}(x;a)=k$ does not have a solution in $x$, then $C_{m-1,a,\tilde{\Gamma}^{-1}(k;a)}=0$.
This recursion immediately leads to a dynamic programming algorithm similar to Algorithm 3, see Algorithm 4. The initial conditions $C_{1,n,k}$ are either one or zero, depending on whether or not the equations $\tilde{\Gamma}(0;s)=k$ have a solution $1\leqslant s\leqslant n$.
  Input: Rank-ordering function $\tilde{h}$ such that for any $r\geqslant 1$, there is a strictly increasing function $\tilde{\Gamma}(\cdot;r):\{0,1,\ldots\}\to\{0,1,\ldots\}$ such that for any $\kappa\geqslant 1$, $\tilde{h}(r_{1:\kappa})=\tilde{\Gamma}(\tilde{h}(r_{1:(\kappa-1)});r_{\kappa})$. Number of summands $m$, maximum summand $n$, target sum $k$
  Initialize $C_{1,\tilde{n},\tilde{k}}=I(1\leqslant\tilde{k}\leqslant\tilde{n})$ for $\tilde{n}\in[n],\tilde{k}\in[k]$
  for $\tilde{m}=2$ to $m$ do
     for $\tilde{k}=1$ to $k$ do
        for $\tilde{n}=1$ to $n$ do
           $C_{\tilde{m},\tilde{n},\tilde{k}}\leftarrow\sum_{a=1}^{\tilde{n}}C_{\tilde{m}-1,a,\tilde{\Gamma}^{-1}(\tilde{k};a)}$
        end for
     end for
  end for
  Output: $C_{m,n,k}$
Algorithm 4 Computation of $C_{m,n,k}$ for a general compositional rank-ordering function $\tilde{h}$
Computation of endpoints. The computation of the interval endpoints $B_{L},B_{U}$ from (7) can be performed efficiently in a similar way. For concreteness, we consider $B_{U}$, and the reasoning for $B_{L}$ is entirely analogous. Now, $B_{U}=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\leqslant q_{U}\right\}.$
For illustration, we will again first consider the case where $h(S_{(r_{1})},\ldots,S_{(r_{m})})=S_{(r_{1})}+S_{(r_{2})}+\ldots+S_{(r_{m})}$ and $\tilde{h}(r_{1:m})=r_{1}+\ldots+r_{m}$ for all $r_{1:m}$, $S_{(r_{1})},\ldots,S_{(r_{m})}$. The problem becomes to compute
$M_{m,n,q}:=M_{m,n,q}(S_{1},\ldots,S_{n}):=\max\{S_{(r_{1})}+S_{(r_{2})}+\ldots+S_{(r_{m})}\mid r_{1}+\ldots+r_{m}\leqslant q\}.$
As above, we can obtain a recursion by considering the possible values of $r_{m}$, to find that $M_{m,n,q}=\max\{M_{m-1,a,q-a}\mid 1\leqslant a\leqslant\min(n,q-m+1)\}$. This recursion can be initialized with $M_{1,n,q}=S_{\min(n,q)}$, leading to a similar dynamic programming algorithm (Algorithm 5).
  Input: number of summands $m$, maximum summand $n$, upper bound on sum $q$
  Initialize $M_{1,\tilde{n},\tilde{k}}=S_{\min(\tilde{n},\tilde{q})}$ for $\tilde{n}\in[n],\tilde{q}\in[q]$
  for $\tilde{m}=2$ to $m$ do
     for $\tilde{q}=1$ to $q$ do
        for $\tilde{n}=1$ to $n$ do
           $M_{\tilde{m},\tilde{n},\tilde{q}}=\max\{M_{\tilde{m}-1,a,\tilde{q}-a}\mid 1\leqslant a\leqslant\min(\tilde{n},\tilde{q}-\tilde{m}+1)\}$
        end for
     end for
  end for
  Output: $M_{m,n,q}$
Algorithm 5 Computation of $M_{m,n,q}$
More generally, consider a finite set $\mathcal{H}\subset\mathbb{R}$ containing, for all $\kappa\in[m]$, the values $h(S_{(r_{1})},\ldots,S_{(r_{\kappa})})$ $1\leqslant r_{1}\leqslant\ldots\leqslant r_{\kappa}\leqslant n$. Suppose that (12) holds, and that similarly, for all $r\geqslant 1$, there is a strictly increasing function $\Gamma(\cdot;r):\mathcal{H}\to\mathcal{H}$ such that for any $\kappa\geqslant 1$,
$h(S_{(r_{1})},\ldots,S_{(r_{\kappa})})=\Gamma(h(S_{(r_{1})},\ldots,S_{(r_{\kappa-1})});r_{\kappa}).$
For instance, for $h(S_{(r_{1})},\ldots,S_{(r_{m})})=S_{(r_{1})}+S_{(r_{2})}+\ldots+S_{(r_{m})}$, we have $\Gamma(a;r)=a+S_{(r)}$. Denote $M_{m,n,q}=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\leqslant q\right\}.$ Then, as above, we can obtain a recursion by considering the possible values of $r_{m}$, to find that $M_{m,n,q}=\max\{\Gamma(M_{m-1,a,\tilde{\Gamma}^{-1}(q;a)};a)\mid 1\leqslant a\leqslant n\}$. By setting the initial conditions $M_{1,n,q}=h(S_{(\tilde{\Gamma}^{-1}(q;n))}),$ we can obtain a dynamic programming algorithm similar to the ones presented above for efficiently computing $M_{m,n,q}$.
Remark 2.
If the calibration and test set sizes are very large, the above algorithms can still have a high cost. However, in certain cases of interest, especially for the central case of the mean, a straightforward procedure for inference is based on concentration inequalities. For instance, if $Y\in[a,b]$ almost surely, then by Hoeffding’s inequality, the prediction set
$\widehat{C}(\mathcal{D}_{n})=\left(\frac{1}{n}\sum_{i=1}^{n}Y_{i}\pm(b-a)\sqrt{\frac{1}{2}\left(\frac{1}{n}+\frac{1}{m}\right)\log\frac{2}{\alpha}}\right)\cap[a,b]$
has $(1-\alpha)$ coverage for the mean of test outcomes. Thus, very large sample sizes can be handled with concentration inequalities, while for moderate sample sizes, our algorithms remain computationally efficient, whereas the concentration-based method may result in trivial prediction sets.
2.3 Simultaneous inference on multiple quantiles
So far, we have proposed methods for univariate target functions $g$ or $h$. In this section, we extend the idea of batch PI to provide a simultaneous prediction set for multiple quantiles of the scores, e.g., $h(s_{1},\cdots,s_{m})=(s_{\zeta_{1}},s_{\zeta_{2}})^{\top}$ for all $s_{1:m}$. This will allow us to provide fine-grained control of the test distribution, for instance by obtaining a prediction set for the interquartile range.
Specifically, we examine the problem of constructing simultaneous bounds for multiple quantiles of test scores. Suppose the target function is given as $h:(s_{1},\cdots,s_{m})\mapsto(s_{(t_{1})},\cdots,s_{(t_{l})})^{\top}$, where $1\leqslant t_{1}\leqslant\cdots\leqslant t_{l}\leqslant m$, and we aim to construct vectors $L=(L_{1},\cdots,L_{l})^{\top}$ and $U=(U_{1},\cdots,U_{l})^{\top}$ serving as bounds such that
$\mathbb{P}\left\{{L\preceq h(S_{(n+1)},\cdots,S_{(n+m)})\preceq U}\right\}=\mathbb{P}\left\{{L_{1}\leqslant S_{(t_{1})}^{\textnormal{test}}\leqslant U_{1},\cdots,L_{n}\leqslant S_{(t_{l})}^{\textnormal{test}}\leqslant U_{l}}\right\}\geqslant 1-\alpha.$ (13)
To provide a procedure that attains the above guarantee, we first introduce some notation. For any $11\leqslant\rho_{1}\leqslant\ldots\leqslant\rho_{l}\leqslant n+$ 1, we will need to compute the number of solutions $r_{1:m}\in H$ of $r_{t_{1}}=\rho_{1},\ldots,r_{t_{l}}=\rho_{l}$. This equals
$\begin{split}L(\rho_{1},\ldots,\rho_{l})&:=\left|\left\{(r_{1},\ldots,r_{m})\in H:r_{t_{1}}=\rho_{1},\ldots,r_{t_{l}}=\rho_{l}\right\}\right|\\ &=_{\rho_{1}}\mathrm{H}_{t_{1}-1}\cdot\left[\prod_{j=1}^{n}{}_{\rho_{j+1}-\rho_{j}+1}\mathrm{H}_{t_{j+1}-t_{j}-1}\right]\cdot_{n-\rho_{l}+2}\mathrm{H}_{m-t_{l}}.\end{split}$ (14)
Next, define for $(w_{1},w_{2},\cdots,w_{l}),(q_{1},q_{2},\cdots,q_{l})$ satisfying $11\leqslant w_{j}\leqslant q_{j}\leqslant n+$ 1 for all $j\in[n+1]$,
$F_{n,m}(w_{1},w_{2},\cdots,w_{l};q_{1},q_{2},\cdots,q_{l})$ =|$=\left|\left\{(r_{1},r_{2},\cdots,r_{m})\in H,w_{j}\leqslant r_{t_{j}}\leqslant q_{j},\;\forall\;j\in[l]\right\}\right|$
$=\sum_{\rho_{1}=w_{1}}^{q_{1}}\sum_{\rho_{2}=\max\{\rho_{1},w_{2}\}}^{q_{2}}\cdots\sum_{\rho_{m}=\max\{\rho_{m-1},w_{l}\}}^{q_{l}}L(\rho_{1},\cdots,\rho_{l}).$ (15)
Applying the idea from the proof of batch PI, we can derive the following result.
Theorem 2.
Suppose that the data points $Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}$ are exchangeable, and that $(w_{1},w_{2},\cdots,w_{l})$ and $(q_{1},q_{2},\cdots,q_{l})$ satisfy $F_{n,m}(w_{1},w_{2},\cdots,w_{l};q_{1},q_{2},\cdots,q_{l})\geqslant(1-\alpha)\cdot\binom{n+m}{m}.$ Let $S_{(0)}=\inf s$ and $S_{(n+1)}=\sup s$. Then
$\mathbb{P}\left\{{S_{(w_{1}-1)}\leqslant S_{(t_{1})}^{\textnormal{test}}\leqslant S_{(q_{1})},S_{(w_{2}-1)}\leqslant S_{(t_{2})}^{\textnormal{test}}\leqslant S_{(q_{2})},\cdots,S_{(w_{l}-1)}\leqslant S_{(t_{l})}^{\textnormal{test}}\leqslant S_{(q_{l})}}\right\}\geqslant 1-\alpha.$
We also mention that Gazin et al. (2024) provided an approach that they refer to as ‘template’, which could also be used to derive joint prediction sets for the order statistics of the test scores.
Continuing with our argument, it remains to determine vectors $(w_{1},\cdots,w_{l})$ and $(q_{1},\cdots,q_{l})$ that satisfy the condition of Theorem 2. For instance, we can consider the following procedure. Let $\tilde{t}_{j}=\text{round}(t_{j}\cdot n/m)$ for $j\in[l]$ represent—roughly speaking—the expected rank of the $j$-th largest test score among the $n$ calibration scores. Then our idea is to center the indices $w_{j}=\tilde{t}_{j}-a$, $q_{j}=\tilde{t}_{j}+a$, $a\geqslant 0$, around $\tilde{t}_{j}$, for $j\in[l]$. Then, we find the smallest $a\in\mathbb{N}$ such that
$F_{n,m}\left((\tilde{t}_{1}-a)\vee 1,\cdots,(\tilde{t}_{l}-a)\vee 1;(\tilde{t}_{1}+a)\wedge(n+1),\cdots,(\tilde{t}_{l}+a)\wedge(n+1)\right)\\ \geqslant(1-\alpha)\binom{n+m}{m},$
and denote it by $t$. Then define
$L=(S_{((\tilde{t}_{1}-t-1)_{+})},\cdots,S_{((\tilde{t}_{2}-t-1)_{+})}),\quad U=(S_{(\min\{\tilde{t}_{l}+t,n+1\})},S_{(\min\{\tilde{t}_{2}+t,n+1\})},\cdots,S_{(\min\{\tilde{t}_{l}+t,n+1\})}).$ (16)
Applying Theorem 2, we have the following result.
Corollary 2.
Suppose the data points $Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}$ are exchangeable. Then for $L$ and $U$ defined in (16), it holds that $\mathbb{P}\left\{{L\preceq(S_{(t_{1})}^{\textnormal{test}},S_{(t_{2})}^{\textnormal{test}},\cdots,S_{(t_{l})}^{\textnormal{test}})\preceq U}\right\}\geqslant 1-\alpha$.
In Section 4.3, we provide experimental results for the specific case of inference on quartiles $S_{(\text{round}(0.25m))}^{\textnormal{test}},S_{(\text{round}(0.75m))}^{\textnormal{test}}$ with the following guarantee: $\mathbb{P}\left\{{L\leqslant S_{(\text{round}(0.25m))}^{\textnormal{test}}\leqslant S_{(\text{round}(0.75m))}^{\textnormal{test}}\leqslant U}\right\}\geqslant 1-\alpha$. For clarity, we include the specific procedure for this task below.
Input: Calibration data $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$. Score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. Test set size $m$. Target coverage level $1-\alpha\in[0,1]$.
Step 1 Compute $t_{1}=\text{round}(0.25\cdot m)$, $t_{2}=\text{round}(0.75\cdot m)$, $\tilde{t}_{1}=\text{round}(0.25\cdot n)$ and $\tilde{t}_{2}=\text{round}(0.75\cdot n)$.
Step 2: Compute $t=\min\left\{a\in\mathbb{N}:\sum\limits_{\rho_{1}=\max\{\tilde{t}_{1}-a,1\}}^{\min\{\tilde{t}_{2}+a,n+1\}}\sum\limits_{\rho_{2}=\rho_{1}}^{\min\{\tilde{t}_{2}+a,n+1\}}{}_{\rho_{1}}\mathrm{H}_{t_{1}-1}\cdot{}_{\rho_{2}-\rho_{1}+1}\mathrm{H}_{t_{2}-t_{1}-1}\cdot_{n-\rho_{2}+2}\mathrm{H}_{m-t_{2}}\geqslant(1-\alpha)\cdot\binom{n+m}{m}\right\}$.
Step 3: Compute the scores $S_{i}=s(X_{i},Y_{i})$ for $i=1,2,\ldots,n$; and let $S_{(0)}=\inf s$ and $S_{(n+1)}=\sup s$.
Return: Bounds $L=S_{(\max\{\tilde{t}_{1}-t-1,0\})}$ and $U=S_{(\min\{\tilde{t}_{2}+t,n+1\})})$.
Algorithm 6 Batch Predictive Inference for quartiles
This procedure also leads to valid inference on the interquartile range $\text{IQR}=S_{(\text{round}(0.75m))}^{\text{test}}-S_{(\text{round}(0.25m))}^{\text{test}}$, with the guarantee $\mathbb{P}\left\{{\text{IQR}\leqslant U-L}\right\}\geqslant 1-\alpha$.
2.4 Inference under covariate shift
Our methods presented so far are valid when the test and calibration data are drawn from the same population, but this might not always hold in applications. This phenomenon has been referred to as dataset shift (see, e.g., Quiñonero-Candela et al., 2009; Shimodaira, 2000; Sugiyama and Kawanabe, 2012). An important form of dataset shift is covariate shift: a changed feature distribution, and an unchanged distribution of the outcome given features. The shift may arise due to a change in the sampling probabilities of various sub-populations, or due to a patient’s features changing over time, while the distribution of the outcome given the features stays fixed (Quiñonero-Candela et al., 2009). There is a growing body of work on distribution-free predictive inference under covariate shift, see e.g., Tibshirani et al. (2019); Qiu et al. (2023); Yang et al. (2023+); Park et al. (2022a); Cauchois et al. (2024); Lei and Candès (2021). However, to our knowledge, methods for batch predictive inference have not been developed yet in this setting.
In this section, we develop methods for batch predictive inference under covariate shift. This refers to the following distribution of the data points:
$\begin{split}&(X_{1},Y_{1}),(X_{2},Y_{2})\ldots,(X_{n},Y_{n})\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X}\times P_{Y\mid X},\\ &(X_{n+1},Y_{n+1}),(X_{n+2},Y_{n+2})\ldots,(X_{n+m},Y_{n+m})\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}Q_{X}\times P_{Y\mid X},\end{split}$ (17)
where $P_{X}$ and $Q_{X}$ represent two distinct distributions on $\mathcal{X}$, and $P_{Y\mid X}$ denotes the conditional distribution of $Y$ given $X$, which is consistent across both the calibration and test datasets. Our objective is to construct a prediction set for a function of the test points under this setting, with coverage at least $1-\alpha$:
$\mathbb{P}_{{Z_{1:n}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X}\times P_{Y\mid X},Z_{(n+1):(n+m)}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}Q_{X}\times P_{Y\mid X}}}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha.$ (18)
Unfortunately, it is known that even in the case of a single test point, constructing a completely distribution-free prediction set under covariate shift is not possible unless the prediction sets are uninformative (cover each fixed value with probability $1-\alpha$)  (Qiu et al., 2023; Yang et al., 2023+). Therefore, as for the case of standard conformal prediction  (Tibshirani et al., 2019), we consider the setting of a known likelihood ratio $dP/dQ$. This setting is of broad interest, arising in randomized trials (e.g., Friedman et al., 2010; Armitage et al., 2013, etc.). and two-phase sampling studies (e.g., Hansen and Hurwitz, 1946; Shrout and Newman, 1989, etc.). As in other problems in conformal prediction, studying the setting where the likelihood ratio is unknown and needs to be estimated might require significant additional development, as well as a different set of tools (Qiu et al., 2023; Yang et al., 2023+).
2.4.1 Reformulation as a missing data problem
To enable a concise argument, it helps to reformulate the problem as a missing data problem. Let $A\in\{0,1\}$ be the binary variable that indicates whether or not the outcome $Y$ is observed. Then the set of all observed data points $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),X_{n+1},X_{n+2},\ldots,X_{n+m}$ can equivalently be viewed as having $n+m$ tuples $(X_{i},A_{i},Y_{i}A_{i})_{1\leqslant i\leqslant n+m}$. The feature distributions $P_{X}$ and $Q_{X}$ in (17) correspond to the conditional distributions $P_{X\mid A=1}$ and $P_{X\mid A=0}$, respectively. Thus, we can rewrite the model (17) as
$\begin{split}&(X_{1},Y_{1}),(X_{2},Y_{2})\ldots,(X_{n},Y_{n})\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X\mid A=1}\times P_{Y\mid X},\\ &(X_{n+1},Y_{n+1}),(X_{n+2},Y_{n+2})\ldots,(X_{n+m},Y_{n+m})\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X\mid A=0}\times P_{Y\mid X},\end{split}$ (19)
and the target coverage guarantee (18) can be written as
$\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\in\widehat{C}(\mathcal{D}_{n})}\ \middle|\ {A_{1},\ldots,A_{n}=1,A_{n+1},\ldots,A_{n+m}=0}\right\}\geqslant 1-\alpha.$
Since the model (19) and the target guarantee do not depend on the marginal distribution of $A$, we are free to assume any value for $\mathbb{P}\left\{{A=1}\right\}$. Note that the tuple ($(\mathbb{P}\left\{{A=1}\right\},P_{X\mid A=1},P_{X\mid A=1})$ determines the joint distribution of $(X,A)$, and thus the distributions $P_{X}$ and $P_{A\mid X}$ are well-defined once $\mathbb{P}\left\{{A=1}\right\}$ is fixed.
From this reframing, knowing the likelihood ratio $dP_{X\mid A=1}/dP_{X\mid A=0}$ can equivalently be thought of as access to the propensity score $x\mapsto p_{A\mid X}(x)=\mathbb{P}\left\{{A=1}\ \middle|\ {X=x}\right\}$ for some value of $\mathbb{P}\left\{{A=1}\right\}$. Indeed, for any $x$,
$\frac{dP_{X\mid A=1}(x)}{dP_{X\mid A=0}(x)}=\frac{\mathbb{P}\left\{{A=1}\ \middle|\ {X}\right\}dP(x)}{\mathbb{P}\left\{{A=0}\ \middle|\ {X}\right\}dP(x)}\cdot\frac{\mathbb{P}\left\{{A=0}\right\}}{\mathbb{P}\left\{{A=1}\right\}}\propto\frac{1-p_{A\mid X}(x)}{p_{A\mid X}(x)}.$
Based on this observation, we start by viewing propensity score as known.
A simple approach one could consider is to extend weighted split conformal prediction. However, as we show in Section A.4, this approach suffers from a similar issue as the standard extension of split conformal prediction. Unless $n\gg m$, it typically results in large prediction sets that can cover the entire range of the random variable of interest.
2.4.2 Proposed method: batch PI with rejection sampling
Input: Calibration data $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$. Propensity score $p_{A\mid X}$ with known pointwise lower bound $c>0$. Score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. Test set size $m$. Batch score function $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$. Rank-ordering function $\tilde{h}:\mathbb{N}^{m}\rightarrow\mathbb{R}$. Target coverage level $1-\alpha\in[0,1]$. Lower and upper error levels $\beta,\gamma\in[0,1]$ satisfying $\beta+\gamma=\alpha$
Step 1: For $i=1,2,\ldots,n$, draw $B_{i}\mid X_{i}\sim\textnormal{Bern}(p_{B\mid X}(X_{i}))$, where $p_{B\mid X}(x)$=$\frac{c}{1-c}\cdot\frac{1-p_{A\mid X}(x)}{p_{A\mid X}(x)}$.
Step 2: Define the subset of the calibration data $\tilde{\mathcal{D}}_{n}=\{(X_{i},Y_{i}):1\leqslant i\leqslant n,B_{i}=1\}$.
Return: Prediction set $\widehat{C}^{\textnormal{bPI-CovShift}}(\mathcal{D}_{n}):=\mathcal{C}^{\textnormal{bPI}}(\tilde{\mathcal{D}}_{n})$, applying batch PI from Algorithm 1 to $\tilde{\mathcal{D}}_{n}$
Algorithm 7 Batch Predictive Inference under Covariate Shift
As an alternative approach, we consider constructing an exchangeable dataset via rejection sampling, as it has been done for standard conformal prediction in Park et al. (2022a); Qiu et al. (2023), and then applying the batch PI procedure.
Suppose we have access to the conditional distribution $P_{A\mid X}$ (again, for some possibly unknown value of $\mathbb{P}\left\{{A=1}\right\}$). We draw a subset of the calibration data set as follows. For each $i=1,2,\ldots,n$, draw
$B_{i}\mid X_{i}\sim\textnormal{Bern}(p_{B\mid X}(X_{i})),\textnormal{ where }p_{B\mid X}(x)=\frac{c}{1-c}\cdot\frac{1-p_{A\mid X}(x)}{p_{A\mid X}(x)}.$ (20)
The Bernoulli distribution described above is well-defined for any value of $X_{i}$ if $p_{A\mid X}(x)>0$ for all $x\in\mathcal{X}$. In fact, for our coverage result, we will require a slightly stronger condition:
Condition 2.
There exists a constant $c\in(0,1)$ such that $p_{A\mid X}(x)\geqslant c$ for all $x\in\mathcal{X}$.
This sampling scheme was previously discussed in Park et al. (2022a), and intuitively, it constructs a subset of the calibration set that mimics the distribution of the test set through reweighting based on the propensity score. Let $\widetilde{\mathcal{D}}_{n}$ be the subset of the calibration data defined as
$\widetilde{\mathcal{D}}_{n}=\{(X_{i},Y_{i}):1\leqslant i\leqslant n,B_{i}=1\}.$ (21)
The subset $\widetilde{\mathcal{D}}_{n}$ of the calibration data is exchangeable with the test data, and thus it follows that the batch PI prediction set $\widehat{C}^{\textnormal{bPI-CovShift}}(\mathcal{D}_{n}):=\widehat{C}(\widetilde{\mathcal{D}}_{n})$ from this subset achieves the target level of coverage:
Corollary 3.
Under Conditions 1 and 2, with $\widetilde{\mathcal{D}}_{n}$ constructed by (21), the batch PI prediction set $\widehat{C}^{\textnormal{bPI-CovShift}}(\mathcal{D}_{n}):=\widehat{C}(\widetilde{\mathcal{D}}_{n})$ based on (8) satisfies $\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\in\widehat{C}(\widetilde{\mathcal{D}}_{n})}\ \middle|\ {A_{1:n},B_{1:n}}\right\}\geqslant 1-\alpha$, where the probability is taken with respect to the model (19).
Similarly, we can combine rejection sampling and the procedure (6) to conduct inference on multiple quantiles of test scores under covariate shift. In general, rejection sampling translates any procedure designed for i.i.d. data to a procedure suitable for data with covariate shift. The procedure $\widehat{C}(\widetilde{\mathcal{D}}_{n})$ is an application of this approach to batch PI. Since rejection sampling reduces the sample size, using naive procedures such as split conformal prediction may yield uninformative prediction sets after rejection sampling, even if the original calibration set is large. The batch PI procedure addresses this issue as its usefulness does not depend heavily on the ratio of calibration to test sizes.
3 Use cases
In this section, we discuss use cases of batch PI: (1) simultaneous predictive inference with dataset-conditional coverage; (2) selection of individuals with error control –—both based on inference on one quantile; and (3) inference on counterfactual variables. All three will be illustrated empirically in Section 4.
3.1 Simultaneous predictive inference of multiple unobserved responses
Consider constructing prediction sets $\widehat{C}_{n}(X_{n+1}),\widehat{C}_{n}(X_{n+2}),\ldots,\widehat{C}_{n}(X_{n+m})$ for $Y_{n+1},Y_{n+2},\ldots,Y_{n+m}$ respectively, such that most of the unobserved outcomes are covered by their corresponding prediction sets. A simple approach is to construct standard split conformal prediction sets, leading to marginal coverage for each prediction set, i.e., $\mathbb{P}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\alpha,\text{ for all }j=1,2,\ldots,m$.
However, this does not characterize the simultaneous—joint—behavior of the prediction sets. For instance, it does not directly guarantee how many of the test outcomes will be covered. Since each marginal coverage guarantee is with respect to the distribution of $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+j},Y_{n+j})$, the $m$ coverage events $\{\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\},j=1,2,\ldots,m\}$ have a joint distribution with a potentially complex dependence structure.
The distribution of the coverage $\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}$ was determined by Marques F. (2025); Huang et al. (2024), and this enables constructing prediction sets various guarantees. Our goal is to construct prediction sets with the following probably approximately correct (PAC)-type (Park et al., 2020) guarantees:
$\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\delta}\right\}\geqslant 1-\alpha,$ (22)
where $\alpha,\delta\in(0,1)$ are predefined levels. This directly controls the proportion of test outcomes covered by the prediction sets. For illustration purposes, we will show that the batch PI procedure can be applied to achieve the above guarantee.
Let $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ be a nonconformity score, constructed independently of the calibration data. Define $m_{\delta}=\lceil(1-\delta)m\rceil$, and the following prediction set, which is a direct application of the procedure for inference on a single quantile:
$\widehat{C}_{n}(x)=\left\{y\in\mathcal{Y}:s(x,y)\leqslant Q_{1-\alpha}\left(\sum_{k=1}^{n+1}\frac{\binom{k+m_{\delta}-2}{m_{\delta}-1}\binom{n+m-k-m_{\delta}+1}{m-m_{\delta}}}{\binom{n+m}{m}}\cdot\delta_{k}\right)\right\}.$ (23)
As a consequence of our general results, we find that if $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),$ $(X_{n+1},Y_{n+1}),$ $\ldots$,… , $(X_{n+m},Y_{n+m})$ are exchangeable, then the prediction set $\widehat{C}_{n}$ from (23) satisfies the guarantee (22). This follows directly by applying the result in Section 2.2.1.
Remark 3 (Comparison with the PAC guarantee for calibration-conditional coverage).
Consider the setting where the data points are i.i.d. Let $C_{j}={\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}$ denote the coverage indicator for the $j$th test point, $j\in[m]$, and let $\mathcal{D}_{\text{cal}}$ denote the calibration set. Then we have $C_{1},\cdots,C_{m}\mid\mathcal{D}_{\text{cal}}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}\textnormal{Bernoulli}(p_{C})$, where $p_{C}=\mathbb{P}\left\{{Y\in\widehat{C}_{n}(X)}\ \middle|\ {\mathcal{D}_{\text{cal}}}\right\}$, and thus $\bar{C}=\frac{1}{m}\sum_{j=1}^{m}C_{j}$ converges to $p_{C}$ almost surely, conditional on $\mathcal{D}_{\text{cal}}$. It follows that
$\mathbb{P}\left\{{\bar{C}\geqslant 1-\delta}\right\}=\mathbb{E}\left[{\mathbb{E}\left[{{\mathbbm{1}}\left\{{\bar{C}\geqslant 1-\delta}\right\}}\ \middle|\ {\mathcal{D}_{\text{cal}}}\right]}\right]\xrightarrow{m\rightarrow\infty}\mathbb{E}\left[{{\mathbbm{1}}\left\{{p_{C}\geqslant 1-\delta}\right\}}\right]=\mathbb{P}\left\{{p_{C}\geqslant 1-\delta}\right\},$
by applying the dominated convergence theorem twice. Therefore, as $m\to\infty$, the prediction set (23) can also be viewed as achieving the PAC guarantee for the calibration conditional-coverage property (Vovk, 2013b; Park et al., 2020) $\mathbb{P}\left\{{p_{C}\geqslant 1-\delta}\right\}\geqslant 1-\alpha$. The advantage of the prediction set (23) is that it controls the coverage rate also for small test sizes $m$.
Remark 4.
This problem was also studied previously in Gazin et al. (2024), where the authors further aim to provide uniform control over the false coverage rate. This can be expressed in our notation as:
$\mathbb{P}\left\{{\forall\,\alpha\in(0,1),\;\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}^{(\alpha)}(X_{n+j})}\right\}\geqslant 1-\gamma_{\alpha,\delta}}\right\}\geqslant 1-\delta,$
where $(\gamma_{\alpha,\delta})_{0<\alpha<1}$ is a family of (random) bounds. They provide a concentration inequality-based approach to achieve this stronger notion of coverage, with a score of the form $S_{i}=|Y_{i}-\hat{\mu}(X_{i},\mathcal{D}_{\text{train}},X_{1:n+m})|$—i.e., the score is constructed using the training data, as well as the calibration and test covariates. They also briefly mention the weaker target (equivalent to (22)) in the appendix, providing method based on an implicit formula. It turns out that explicitly restating their approach is equivalent to the method in (23). In other words, applying the batch PI methodology to false coverage control recovers the approach of Gazin et al. (2024) in a simpler explicit form.
3.2 Selection of test individuals
Next, we consider selecting the individuals in the test set whose outcome values satisfy a certain condition—for instance, selecting individuals whose outcome values exceed a threshold, i.e., $Y_{i}>c$ for some $c\in\mathbb{R}$. This setting was investigated by Jin and Candès (2023b) and Jin and Candès (2023a), where they discuss applications to candidate screening, drug discovery, etc. Denoting the “null” events as $E_{j}=\{Y_{n+j}\leqslant c\},j=1,2,\ldots,m$, we can view this problem as controlling an error measure depending on the number of true events declared to be false. Previous work (Jin and Candès, 2023b, a) has developed methods for controlling a quantity analogous to the false discovery rate (Benjamini and Hochberg, 1995). Here, we introduce a different procedure, which applies batch PI, directly controlling the number of false claims on the test set.
We assume that $Y$ is bounded below—without loss of generality, suppose $Y\geqslant 0$ almost surely. Generally, for unbounded $Y$, we can apply a monotone transformation to obtain a bounded outcome $\tilde{Y}$—e.g., $\tilde{Y}=\tanh(Y)$—and then apply the procedure below. Let $\hat{\mu}:\mathcal{X}\rightarrow\mathbb{R}_{\geqslant 0}$ be an estimated mean function, constructed on a separate independent dataset. Let $s(x,y)=\hat{\mu}(x){\mathbbm{1}}\left\{{y\leqslant c}\right\}$ for all $x,y$, and define $S_{i}=s(X_{i},Y_{i})$ for $i=1,2,\ldots,n$. We write $S_{(1)},\ldots,S_{(n)}$ to denote the order statistics of $S_{1},\ldots,S_{n}$. Next, for a target number of errors $\eta\in\{0\}\cup[m]$, let
$\hat{T}=S_{(q_{\eta})},\text{ where }q_{\eta}=Q_{1-\alpha}\left(\sum_{k=1}^{n+1}\frac{\binom{k+m-\eta-2}{m-\eta-1}\binom{n+\eta-k+1}{\eta}}{\binom{n+m}{m}}\cdot\delta_{k}\right),$
following the formula in (9) with $\zeta=m-\eta$ and $\gamma=\alpha$. Then we consider the following selection rule:
declare Ej to be false if $\textnormal{declare $E_{j}$ to be false if $\hat{\mu}(X_{n+j})>\hat{T}$}.$ (24)
This satisfies the following property:
Corollary 4.
Suppose $\hat{\mu}(X)\geqslant 0$ holds almost surely. Then the selection procedure (24) controls the number of false claims by $\eta$ with probability at least $1-\alpha$, i.e.,
$\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})>\hat{T},Y_{n+j}\leqslant c}\right\}\leqslant\eta}\right\}\geqslant 1-\alpha.$ (25)
If $\eta=0$, then (25) is equivalent to making at least one false claims with probability at most $\alpha$, $\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})>\hat{T},Y_{n+j}\leqslant c}\right\}>0}\right\}$; which is analogous to the control of the family-wise error rate (FWER) in multiple hypothesis testing. More generally, (25) is analogous to the control of the $k$-family-wise error rate ($k$-FWER) (Lehmann and Romano, 2005) in multiple hypothesis testing.
As a remark, if we are generally interested in selecting individuals whose outcome satisfies a condition $\mathcal{C}$ using an estimator $\hat{f}(\cdot)$ (which is nonnegative), we can apply the same procedure with the score function $s(x,y)=\hat{f}(x){\mathbbm{1}}\left\{{y\text{ satisfies }\mathcal{C}}\right\}$, and then select the individuals whose $\hat{f}$ value exceeds $\hat{T}$.
3.3 Inference on counterfactual variables
Finally, we consider a randomized trial setting where the underlying data structure is $(X_{i},A_{i},Y_{i}^{a=0},Y_{i}^{a=1})_{1\leqslant i\leqslant n}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X}\times P_{A\mid X}\times P_{Y^{a=0}\mid X}\times P_{Y^{a=1}\mid X}$, where $X$ denotes the feature, $A\in\{0,1\}$ denotes the treatment, and $Y^{a=0}$ and $Y^{a=1}$ denote the counterfactual outcomes under $A=0$ and $A=1$, respectively. We only observe $(X_{i},A_{i},Y_{i})_{1\leqslant i\leqslant n}$, where we assume the consistency condition $Y_{i}=(1-A_{i})Y_{i}^{a=0}+A_{i}Y_{i}^{a=1}$.
We consider the task of inference on the counterfactual outcomes $\{Y_{i}^{a=0}:A_{i}=1\}$ in the treated group. Under the consistency assumption $Y=(1-A)Y^{a=0}+AY^{a=1}$, the problem is equivalent to inference on missing outcomes/test points under covariate shift, with data points $(X_{i},A_{i},A_{i}Y_{i}^{a=0})$. Indeed, we can regard $\{(X_{i},Y_{i}^{a=0}):A_{i}=0\}$ as the calibration set and $\{X_{i}:A_{i}=1\}$ as the test inputs.
Therefore, based on the discussion in Section 2.4, e.g., by applying the procedure (7), we obtain procedures for the following tasks:
  1. 1.
    Inference on the mean of counterfactuals: Construct $\widehat{C}(\mathcal{D}_{n})$ such that $\mathbb{P}\left\{{\frac{1}{N^{1}}\sum_{i:A_{i}=1}Y_{i}^{a=0}\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$, where $N^{1}=|\{i:A_{i}=1\}|$.
  2. 2.
    Inference on the median of counterfactuals: Construct $\widehat{C}(\mathcal{D}_{n})$ such that $\mathbb{P}\left\{{\textnormal{Median}(\{Y_{i}^{a=0}:A_{i}=1\})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$.
  3. 3.
    Inference on multiple quantiles of counterfactuals: Construct $L,U\in\mathbb{R}^{l}$ such that $\mathbb{P}\left\{{L\preceq(Y_{(\zeta_{1})}^{a=0},\cdots,Y_{(\zeta_{l})}^{a=0})\preceq U}\right\}\geqslant 1-\alpha$, where $Y_{(\zeta)}^{a=0}$ denotes the $\zeta$-th smallest value of $\{Y_{i}^{a=0}:A_{i}=1\}$.
4 Simulations
In this section, we illustrate the performance of batch PI-based procedures across different experiments [^4] Code to reproduce the experiments is available at https://github.com/yhoon31/batch-PI..
4.1 Simultaneous predictive inference of multiple unobserved outcomes
We generate the data according to the distribution $X\sim N_{p}(\mu_{x},5\cdot I_{p}),Y\mid X\sim\mathcal{N}(\beta_{1}^{\top}X+(\beta_{2}^{\top}X)^{2},|\beta_{3}^{\top}X|^{2})$, where we set the dimension as $p=20$, and the mean vectors $\mu_{x}$ and $\beta_{1},\beta_{2},\beta_{3}$ are randomly generated by drawing each component from uniform distributions over the unit interval. First, we generate a training dataset of size $n_{\text{train}}=200$, and then fit a random forest regression estimator to estimate the mean function $\hat{\mu}(\cdot)$.
Next, we repeat the following steps 500 times: We generate a calibration set of size $n=200$ and a test set of size $m=100$. We then apply the batch PI procedure described in Section 3.1 at level $\delta=0.1$ and $\alpha=0.1,0.05,0.01$. For comparison, we also run split conformal prediction at level $0.1$. The two methods provide the following guarantees, respectively:
[…]
where $\hat{r}=\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}$ denotes the coverage rate over the test set. We sample $\hat{r}$ 500 times for both methods, and compare the estimated means and the probability of $\hat{r}$ exceeding $0.9$. The results are summarized in Table 1 and Figure 2.
$\mathbb{E}\left[{\text{coverage}}\right]$$\mathbb{P}\left\{{\text{coverage}\geqslant 0.9}\right\}$
split conformal 0.9022 (0.0016) 0.6100 (0.0218)
batch PI ($\alpha=0.1$) 0.9366 (0.0012) 0.9280 (0.0116)
batch PI ($\alpha=0.05$) 0.9468 (0.0012) 0.9660 (0.0081)
batch PI ($\alpha=0.01$) 0.9663 (0.0010) 0.9940 (0.0035)
Table 1: Mean of test coverage, probability of test coverage being larger than $0.9$, and the mean prediction interval width of the split conformal and batch PI prediction sets, with standard errors.
Figure 2: Test coverage rates and prediction interval widths of split conformal and batch PI prediction sets.
Table 1 shows that both methods achieve their target guarantees tightly. As further supported by Figure 2, the batch PI-based method achieves stronger control over the test coverage rate by permitting slightly wider prediction sets. Specifically, in all three settings ($\alpha=0.1,0.05,0.01$), the test coverage rate of batch PI exceeds 0.9 in a fraction $(1-\alpha)$ of the trials. In contrast, the split conformal method, aimed at controlling the marginal coverage rate, allows the test coverage rate to fall below 0.9 in many of the trials, while providing a shorter prediction set. The second plot of Figure 2 illustrates this tradeoff between the width of the prediction set and the strength of the target guarantee.
4.2 Selection with error control
Next, we illustrate the performance of batch PI procedure for the selection task described in 3.2. We generate the data from the distribution $X\sim N_{p}(\mu_{x},5\cdot I_{p}),Y=\log(1+\exp(\beta^{\top}X+\sigma Z)),\text{ where }Z\sim\mathcal{N}(0,1)$. The dimension is set to $p=20$, $\sigma=3$, and the mean vectors $\mu_{x}$ and $\beta$ are generated by drawing each component from uniform distributions over the unit interval. We consider the task of selecting individuals with $Y>5$, while controlling the number of false claims, i.e., the number of individuals selected whose actual outcome is five or less.
We first generate a training data of size $n_{\text{train}}=500$, and then fit a random forest regression to construct the score function $s:(x,y)\mapsto\hat{\mu}(x){\mathbbm{1}}\left\{{y\leqslant 5}\right\}$. Next, we repeat the process of generating calibration data of size $n=1000$ and test data of size $m=100$, 500 times. In each trial, we run the selection procedure (24) at level $\alpha=0.1$ and $0.2$, with $\eta=0,2,4,6,8,10$. We record the number of false claims, as well as the number of true claims in each trial. The results are summarized in Table 2 and Figure 3, illustrating that the proposed procedure controls the number of false claims across various target levels $\eta$, satisfying the guarantee (25).
$\eta=0$ $\eta=2$ $\eta=4$ $\eta=6$ $\eta=8$ $\eta=10$
[…]
{# false $\mathbb{P}\left\{{\text{\# false claims}>\eta}\right\}$ 0.084 (0.0006) 0.092 (0.0006) 0.104 (0.0006) 0.102 (0.0006) 0.122 (0.0007) 0.092 (0.0006)
[…]
{# false $\mathbb{P}\left\{{\text{\# false claims}>\eta}\right\}$ 0.172 (0.0008) 0.204 (0.0008) 0.212 (0.0008) 0.206 (0.0008) 0.206 (0.0008) 0.194 (0.0008)
Table 2: Mean of the number of false claims, probability of the number of false claims being larger than the target level $\eta$, and the power of the batch PI-based selection procedure, with standard errors, for $\eta=0,2,4,6,8,10$ and $\alpha=0.1$.
Figure 3: Number of false claims, probability of the number of false claims being larger than the target level $\eta$, and the power of the batch PI-based selection procedure, under $\eta=0,2,4,6,8,10$ and $\alpha=0.1,0.2$.
4.3 Inference on counterfactual variables
In this section, we provide experimental results for the predictive inference on counterfactual variables. We generate the data as $(X_{i},A_{i},Y_{i}^{a=0},Y_{i}^{a=1})\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X}\times P_{A\mid X}\times P_{Y^{a=0}\mid X}\times P_{Y^{a=1}\mid X}$, where $P_{X}$ is an entry-wise uniform distribution on $[0,1]^{p}$, and the treatment $A$ is assigned based on the logistic model $\mathrm{logit}\,\mathbb{P}\left\{{A=1}\ \middle|\ {X=x}\right\}=\beta_{A}^{\top}x$ for all $x$, where the parameter $\beta_{A}\in\mathbb{R}^{p}$ is generated randomly from a uniform distributions over $[0,1]^{p}$. The counterfactual distributions are set as
$Y^{a=0}\mid X\sim\textnormal{Beta}(1+X^{\top}\beta_{Y},1-X^{\top}\beta_{Y}),\qquad Y^{a=1}\mid X\sim\textnormal{Beta}(1-X^{\top}\beta_{Y},1+X^{\top}\beta_{Y}),$
where the parameter $\beta_{Y}$ is generated randomly from a uniform distribution $[0,1]^{p}$.
We first illustrate the performance of our procedure for inference on the quantiles of counterfactual variables. We conduct experiments with a calibration (untreated group) size of $n=200$ and test (treated group) size of $m=40$—i.e., we investigate treatment-conditional inference where the treatment assignments are given. We consider the following tasks:
  1. 1.
    Inference on the median: Find $L,U$ such that $\mathbb{P}\left\{{L\leqslant Y_{(20)}^{a=0}\leqslant U}\right\}\geqslant 1-\alpha$.
  2. 2.
    Simultaneous inference on quartiles: Find $L,U$ such that $\mathbb{P}\left\{{L\leqslant Y_{(10)}^{a=0}\text{ and }Y_{(30)}^{a=0}\leqslant U}\right\}\geqslant 1-\alpha$.
Here, $Y_{(\zeta)}^{a=0}$ denotes the $\zeta$-th smallest value among $\{Y_{n+j}^{a=0}:j=1,2,\cdots,m\}$.
We repeat the process of generating the calibration and test sets, and then applying the procedures at levels $\alpha=0.05,0.075,0.1,\cdots,0.2$, 500 times; and then compute the coverage rates. The results are summarized in Table 3 and 4, illustrating that the procedure tightly attains the target coverage rate.
Target $\alpha=0.05$ $\alpha=0.075$ $\alpha=0.1$ $\alpha=0.125$ $\alpha=0.15$ $\alpha=0.175$ $\alpha=0.2$
Median 0.970 (0.0076) 0.940 (0.0106) 0.924 (0.0119) 0.906 (0.0131) 0.878 (0.0147) 0.852 (0.0159) 0.834 (0.0167)
Quartiles 0.972 (0.0074) 0.954 (0.0094) 0.932 (0.0113) 0.904 (0.0132) 0.868 (0.0152) 0.840 (0.0164) 0.814 (0.0174)
Table 3: Coverage rates of the batch PI prediction sets for counterfactual quartiles (upper: median, lower: quartiles) at different levels, with standard errors.
Figure 4: Coverage rates of the batch PI prediction sets for the median and the quartiles of counterfactual variables at different levels. The dotted line corresponds to $y=x$ line.
Next, we investigate the task of inference on the mean of counterfactual variables, where we aim to construct a bound $B$ that satisfies $\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}Y_{n+j}^{a=0}\leqslant B}\right\}\geqslant 1-\alpha$. We perform the experiment with a calibration size of $n=100$ and the test sizes of $m=5$ and $m=10$. The calibration size after rejection sampling is smaller—around 40 in this experiment. Thus, neither the naive method (which requires a sufficiently large calibration-to-test ratio) nor the concentration-based method (which requires large calibration and test sizes) is useful—they both provide trivial prediction sets of $[0,1]$.
We repeatedly generate the data and run the batch PI procedure with the dynamic programming approach 4 500 times (which uses the rank-ordering function $\tilde{h}(r_{1},\cdots,r_{m})=\sum_{j=1}^{m}r_{j}$), and compute the coverage rate. The results are shown in Table 4 and Figure 5.
Test size $\alpha=0.05$ $\alpha=0.075$ $\alpha=0.1$ $\alpha=0.125$ $\alpha=0.15$ $\alpha=0.175$ $\alpha=0.2$
$m=5$ 0.976 (0.0069) 0.966 (0.0081) 0.954 (0.0094) 0.938 (0.0108) 0.928 (0.0116) 0.926 (0.0117) 0.916 (0.0124)
$m=10$ 0.982 (0.0060) 0.972 (0.0074) 0.962 (0.0086) 0.954 (0.0094) 0.848 (0.0099) 0.840 (0.0106) 0.824 (0.0119)
Table 4: Coverage rates of the prediction set for the mean of counterfactual variables at different levels, with standard errors.
Figure 5: Coverage rates of the prediction set for the mean of counterfactual variables for test sizes five and ten, at different levels. The dotted line corresponds to $y=x$ line.
The results show that the batch PI prediction set achieves the coverage guarantee, though it is a bit conservative. This conservativeness is partly due to the discrepancy between the the rank-ordering function $\tilde{h}$ and the actual ordering of the $h$ values, which we do not have access to in practice.
To further illustrate this, we empirically examine the coverage rates of the batch PI prediction sets for the mean of the test scores under various score distributions with bounded support, with calibration and test sizes set to $n=40$ and $m=10$, respectively. Figure 6 demonstrates that the batch PI procedure achieves the target coverage guarantee across different distributions, though with varying levels of tightness. While the prediction set is designed to ensure a distribution-free guarantee—controlling for worst-case scenarios—it may be conservative in general. Nonetheless, these prediction sets remain useful in some sense, as neither naive methods nor concentration-based methods provide nontrivial prediction sets in this setting.
Figure 6: Coverage rates of the prediction set for the mean of test scores under various score distributions. The left plot visualizes the score distributions, while the right plot shows the coverage rates of the batch PI prediction sets. The dotted line represents the $y=x$ line.
In Appendix D, we provide simulation results in the setting where we do not have access to the true propensity score and instead use an estimate of the propensity score in the procedure. These results demonstrate that our methodology remains robust, yielding similar results even when relying on the estimates.
5 Empirical data illustration
Next, we illustrate the performance of the batch PI procedure by applying it to a drug-target interaction (DTI) dataset. We use the dataset and the pre-trained model from the DeepPurpose library (Huang et al., 2020). The original dataset has 16,486 observations in both the calibration and the test sets. The covariates consist of a pair of drug and target protein, and the response variable is the affinity score, which is a real-valued measure of the interaction between the drug and the target protein.
We first consider the task of constructing prediction sets for each unobserved outcome variable—as discussed in Section 3.1. To illustrate performance under moderate sample sizes, we create a calibration set of size 500 randomly drawn from the original calibration data. We then construct 160 test sets, each of size 100, using a total of 16,000 observations from the test set. Denoting the pretrained estimator by $\hat{\mu}$, we run the batch PI-based procedure (23) with the score $s:(x,y)\mapsto|y-\hat{\mu}(x)|$ at levels $\delta=0.1$ and $\alpha=0.05,0.1,0.15,0.2,0.25,0.3$. For comparison, we also run split conformal prediction at level $\delta=0.1$ for each of the test points. We compute the proportion of test sets (out of 160 total sets) where the coverage rate exceeds 0.9, as well as the mean coverage rate. The results are summarized in Table 5 and Figure 7.
batch PI ($\delta=0.1$) split conformal $(\text{level}=0.1)$
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.15$ $\alpha=0.2$ $\alpha=0.25$ $\alpha=0.3$
$\mathbb{P}\left\{{\textnormal{coverage}\geqslant 0.9}\right\}$ 0.9500 (0.0173) 0.9312 (0.0201) 0.8438 (0.0288) 0.7938 (0.0321) 0.7812 (0.0328) 0.7500 (0.0343) 0.6062 (0.0387)
$\mathbb{E}\left[{\textnormal{coverage}}\right]$ 0.9406 (0.0019) 0.9350 (0.0021) 0.9249 (0.0023) 0.9178 (0.0024) 0.9152 (0.0024) 0.9131 (0.0024) 0.9040 (0.0024)
Table 5: The proportion of test sets with test-coverage being larger than $0.9$, and the mean coverage rate of the batch PI-based procedure at levels $\alpha=0.05,0.1,\cdots,0.3$ and $\delta=0.1$, along with the split conformal-based procedure at level $0.1$, with standard errors.
Figure 7: The proportion of test sets whose test coverage exceeds 0.9, and the mean coverage rate of the batch PI and split conformal-based procedures at different levels. The dotted lines represent the $y=x$ line (left) and the $y=0.9$ line (right), respectively.
The results illustrate that both methods attain their respective target guarantees. The batch PI-based procedure controls the probability of the test coverage exceeding 0.9 at different values of $\alpha$, whereas the split conformal method does not control this probability, and instead controls the mean coverage rate tightly.
Next, we examine the task of selecting drug-target pairs with high scores, following the discussion in Section 3.2. We construct a calibration set of size 2000, and 160 test sets of size 100. We aim to select drug-protein pairs whose corresponding scores exceed seven, which roughly corresponds to the top 20% of the entire set. We run the procedure (24) at levels $\alpha=0.05,0.1,0.15,0.2,0.25,0.3$ and target numbers of false claims $\eta=0,3,5$ (recall that the procedure at $\eta=0$ controls a quantity analogous to the family-wise error rate (FWER)). The results are shown in Table 6 and Figure 8, illustrating that the batch PI procedure achieves the target guarantee at various levels.
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.15$ $\alpha=0.2$ $\alpha=0.25$ $\alpha=0.3$
$\eta=0$ 0.0437 (0.0162) 0.1250 (0.0262) 0.1938 (0.0313) 0.2125 (0.0324) 0.2250 (0.0331) 0.2375 (0.0337)
{# false $\mathbb{P}\left\{{\textnormal{\# false claims}\geqslant\eta}\right\}$ $\eta=3$ 0.0312 (0.0138) 0.0625 (0.0192) 0.1250 (0.0262) 0.1625 (0.0293) 0.2375 (0.0337) 0.2562 (0.0346)
$\eta=5$ 0.0125 (0.0088) 0.0625 (0.0192) 0.1375 (0.0273) 0.2188 (0.0328) 0.2688 (0.0352) 0.3000 (0.0363)
Table 6: The proportion of test sets whose number of false claims exceeds the target $\eta$, at levels $\alpha=0.05,0.1,\cdots,0.3$ and $\eta=0,3,5$, with standard errors.
Figure 8: The proportion of test sets whose number of false claims exceeds the target $\eta$, at levels $\alpha=0.05,0.1,\cdots,0.3$ and $\eta=0.3,5$. The dotted lines represent the $y=x$ line.
6 Discussion
This work introduces a distribution-free framework for joint predictive inference on a batch of multiple test points. The proposed batch PI method, provides procedures for various inference problems, such as constructing multiple prediction sets with PAC-type guarantees, constructing a selection procedure that controls the number of false claims, and inference on the mean or median of unobserved outcomes.
Many open questions remain. For inference on one test point, several works have explored developing new distribution-free procedures that can achieve stronger targets or operate under more complex data structures. Examples include attaining training- or test-conditional coverage guarantees, or developing methods that work with non-exchangeable data. Similar questions can be asked for joint inference on multiple objects. For example, can we achieve batch-conditional inference, and what kind of conditional coverage can be controlled? If we have a hierarchical structure in the data involving groups of observations, how can we perform inference for new groups?
Another important direction would be to expand the range of targets that the inference framework can address. For example, if one is more interested in the proportion rather than the number of false claims in the selection task, can we construct a procedure with a guarantee of the form, $\mathbb{P}\left\{{\text{FDP}>\alpha}\right\}<\delta$? Can we generally target a functional of the empirical distribution of the unobserved outcomes? We leave these questions to future work.
Acknowledgement
This work was supported in part by NIH R01-AG065276, R01-GM139926, NSF 2210662, P01-AG041710, R01-CA222147, ARO W911NF-23-1-0296, NSF 2046874, ONR N00014-21-1-2843, and the Sloan Foundation. We thank Gilles Blanchard, Ulysse Gazin, and Etienne Roquain for helpful discussion.
Appendix A Naive approaches
A.1 Partitioning the calibration data
A potential approach to achieve (1) is to partition the calibration data, to obtain multiple groups of observations that are exchangeable with the test set. Specifically, suppose $n=mq+r$ where $q$ is a non-negative integer and $10\leqslant r\leqslant m-$ 1. Let $\tilde{Z}_{k}=\{Z_{(k-1)m+1},Z_{(k-1)m+2},\ldots,Z_{km}\}$ for $k\in[q]$ and $\tilde{Z}_{\textnormal{test}}=\{Z_{n+1},\ldots,Z_{n+m}\}$. Then it is clear that $g(\tilde{Z}_{1}),g(\tilde{Z}_{2}),\ldots,g(\tilde{Z}_{q}),g(\tilde{Z}_{\textnormal{test}})$ are exchangeable, and thus we can apply split conformal prediction to obtain the following prediction set for $g(\tilde{Z}_{\textnormal{test}})$:
$\widehat{C}(\mathcal{D}_{n})=\left[Q_{\beta}^{\prime}\left(\sum_{k=1}^{q}\frac{1}{q+1}\delta_{g(\tilde{Z}_{k})}+\frac{1}{q+1}\delta_{\infty}\right),Q_{1-\gamma}\left(\sum_{k=1}^{q}\frac{1}{q+1}\delta_{g(\tilde{Z}_{k})}+\frac{1}{q+1}\delta_{\infty}\right)\right],$ (27)
where $\beta,\gamma\in[0,1]$ satisfies $\beta+\gamma=\alpha$. For example one can set $\beta=\gamma=\alpha/2$ for the construction of a two-sided prediction interval, while $\beta=0,\gamma=\alpha$ yields a one-sided interval. The above method achieves the coverage guarantee (1), but the usefulness is limited to the case where $n\gg m$. For example, if $n<m(1/\alpha-1)$ so that $q+1<1/\alpha$ holds, then it leads to a trivial prediction set.
A.2 Extending split conformal prediction
Instead of constructing exchangeable groups, one can directly leverage individual-level exchangeability. Let
$\begin{split}\overline{S}_{i}&=S_{i}{\mathbbm{1}}\left\{{1\leqslant i\leqslant n}\right\}+(\sup s){\mathbbm{1}}\left\{{n+1\leqslant i\leqslant n+m}\right\},\\ \underline{S}_{i}&=S_{i}{\mathbbm{1}}\left\{{1\leqslant i\leqslant n}\right\}+(\inf s){\mathbbm{1}}\left\{{n+1\leqslant i\leqslant n+m}\right\},\end{split}$ (28)
where $S_{i}=s(Z_{i})$. For $s_{1}\leqslant s_{2}\leqslant\ldots\leqslant s_{m}$, we define $h(s_{1},s_{2},\ldots,s_{m})$ as $\sup h$ if $s_{m}=\sup s$ and $h$ is not well-defined, e.g., $\sup s=+\infty$ and $h(s_{1},\ldots,s_{m})=\sum_{j=1}^{m}s_{i}$. Similarly, we define $h(s_{1},s_{2},\ldots,s_{m})$ as $\inf h$ if $s_{1}=\inf s$ and $h$ is not well-defined; while noting that only one of the two cases can occur below. Then, the naive split conformal prediction set $\widehat{C}(\mathcal{D}_{n})$ is defined as:
$\widehat{C}(\mathcal{D}_{n})=\left[Q_{\beta}^{\prime}\left(\sum_{1\leqslant i_{1}<\ldots<i_{m}\leqslant n+m}\frac{1}{\binom{n+m}{m}}\delta_{h(\underline{S}_{i_{1}},\ldots,\underline{S}_{i_{m}})}\right),Q_{1-\gamma}\left(\sum_{1\leqslant i_{1}<\ldots<i_{m}\leqslant n+m}\frac{1}{\binom{n+m}{m}}\delta_{h(\overline{S}_{i_{1}},\ldots,\overline{S}_{i_{m}})}\right)\right],$ (29)
where $\beta,\gamma\geqslant 0$ are predefined levels satisfying $\beta+\gamma=\alpha$. It can be shown that this is a valid distribution-free prediction set, based on arguments similar to those used in the proof for split conformal prediction. Specifically, under Condition 1, the prediction set $\widehat{C}_{n}$ from (29) satisfies the coverage guarantee (1).
However, this approach still faces limitations unless $n\gg m$. For instance, consider the scenario where $n=m$ and $\sup s=+\infty$. Then half of the $(\bar{S}_{i})_{1\leqslant i\leqslant n+m}$ values are +$\infty$+ ∞, likely leading to a trivial upper bound in (29).
A.3 Extending full conformal prediction
To avoid the issue of having a large mass at $\infty$ or $\infty$- ∞, one may try to construct a full conformal-type prediction set instead of relying on split conformal-type constructions. For example, we can first construct a joint prediction set for $(y_{(n+1):(n+m)})$ as
$\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})\\ =\left\{\tilde{y}=(y_{(n+1):(n+m)}):h(S_{(n+1):(n+m)}^{\tilde{y}})\leqslant Q_{1-\alpha}\left(\sum_{1\leqslant i_{1}<\ldots<i_{m}\leqslant n+m}\frac{1}{\binom{n+m}{m}}\delta_{h(S_{i_{1}}^{\tilde{y}},\ldots,S_{i_{m}}^{\tilde{y}})}\right)\right\},$ (30)
where $S_{i}^{\tilde{y}}=s^{\tilde{y}}(X_{i},Y_{i})$ and $s^{\tilde{y}}$ is the nonconformity score constructed from $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $(X_{n+1},y_{n+1}),\ldots,(X_{n+m},y_{n+m})$—this step is essentially equivalent to the transductive conformal prediction Vovk [2013c]. Then the prediction set for $g(\{Z_{n+1},\ldots,Z_{n+m}\})$ can be constructed as $\widehat{C}(\mathcal{D}_{n})=\left\{g(\{(X_{n+1},y_{n+1}),\ldots,(X_{n+m},y_{n+m})\})):(y_{(n+1):(n+m)})\in\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})\right\}$.
However, this full-conformal type procedure suffers greatly from a heavy computational load. Computing the prediction set (30) requires repeating the computation of scores and quantiles for all tuples $(y_{(n+1):(n+m)})$ in $\mathbb{R}^{m}$. Even if we carry out these steps on a grid, the number of steps increases exponentially with the size of the test set, making this procedure computationally infeasible in most practical scenarios.
A.4 Naive method: extending weighted conformal prediction
A simple approach one could consider for inference under covariate shift in Section 2.4 is to extend weighted conformal prediction. Specifically, suppose the propensity score $p_{A\mid X}$ (corresponding to some possibly unknown value of $\mathbb{P}\left\{{A=1}\right\}$) is known. Then, for each subset $I\subset[n+m]$ of size $|I|=m$, define
$p_{A\mid X}(I)=\frac{\prod_{i\in I}(1-p_{A\mid X}(X_{i}))/p_{A\mid X}(X_{i})}{\sum_{I^{\prime}\subset[n+m],|I^{\prime}|=m}\prod_{i\in I^{\prime}}(1-p_{A\mid X}(X_{i}))/p_{A\mid X}(X_{i})}.$
Also define, for each $I=\{i_{1},i_{2},\ldots,i_{m}\}$ with $1\leqslant i_{1}<i_{2}<\ldots<i_{m}\leqslant n+m$, the vectors $\underline{S}_{I}=(\underline{S}_{i_{1}},\underline{S}_{i_{2}},\ldots,\underline{S}_{i_{m}})$, $\overline{S}_{I}=(\overline{S}_{i_{1}},\overline{S}_{i_{2}},\ldots,\overline{S}_{i_{m}})$, where $\underline{S}_{i}$ and $\overline{S}_{i}$ follow the definition in (28). Then we can construct the prediction set
$\widehat{C}(\mathcal{D}_{n})=\left[Q_{\beta}^{\prime}\left(\sum_{I\subset[n+m],|I|=m}p_{A\mid X}(I)\cdot\delta_{h(\underline{S}_{I})}\right),Q_{1-\gamma}\left(\sum_{I\subset[n+m],|I|=m}p_{A\mid X}(I)\cdot\delta_{h(\overline{S}_{I})}\right)\right].$ (31)
This has the following property:
Proposition 2.
Suppose Condition 1 holds and the data is generated by (19). Then the prediction set from (31) satisfies $\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$, where the probability is taken with respect to the model (19).
The prediction set (31), extending weighted split conformal prediction, suffers from a similar issue as the prediction set (29), which extends split conformal prediction. Unless $n\gg m$, a substantial proportion of $\overline{S}_{i}$s take the value $\sup s$ and $\underline{S}_{i}$s take the value $\inf s$, likely resulting in a prediction set with a non-useful width.
Appendix B Additional details
B.1 One-sided batch PI
Input Calibration data $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$. Score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. Test set size $m$. Batch score function $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$. Rank-ordering function $\tilde{h}:\mathbb{N}^{m}\rightarrow\mathbb{R}$. Target coverage level $1-\alpha\in[0,1]$.
Goal: Construct prediction set for $g(s(X_{n+1},Y_{n+1}),\ldots s(X_{n+m},Y_{n+m}))=$ $h((s(X_{n+1},Y_{n+1}),\ldots s(X_{n+m},Y_{n+m}))_{\uparrow})$.
Step 1: With $H=\left\{r_{1:m}:=(r_{1},\ldots,r_{m})^{\top}:1\leqslant r_{1}\leqslant\ldots\leqslant r_{m}\leqslant n+1\right\}$, compute the sample quantile induced by the rank-ordering function $\tilde{h}$: $q=Q_{1-\alpha}\left(\sum_{r_{1:m}\in H}\delta_{\tilde{h}(r_{1:m})}/\binom{n+m}{m}\right)$.
Step 2: Compute the scores $S_{i}=s(X_{i},Y_{i})$ for $i=1,2,\ldots,n$; and $S_{(n+1)}=\sup s$,.
Step 3: Compute the upper bound $B=\max\left\{h(S_{(r_{1})},\ldots,S_{(r_{m})}):r_{1:m}\in H,\,\tilde{h}(r_{1:m})\leqslant q\right\}$.
Return: Prediction set $\widehat{C}(\mathcal{D}_{n})=\big{(}-\infty,B\big{]}$.
Algorithm 8 One-sided Batch Predictive Inference (batch PI)
Appendix C Batch predictive inference for general sparse functions
Here, we describe the simplification of the batch PI procedure for general sparse function targets. As usual, we consider a target function $g$ that satisfies Condition 1, i.e., there exists a monotone function $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$ such that $g(\{z_{1},\ldots,z_{m}\})=h(s(z)_{\uparrow})$. Further, we consider the case where the function $h$ is sparse, meaning there exists a small subset $\{t_{1},\ldots,t_{l}\}\subset[m]$, $t_{1}<\ldots<t_{l}$, such that $h(s_{1},\ldots,s_{m})$ depends only on $(s_{t_{1}},\ldots,s_{t_{l}}$). In other words, there exists a function $h^{\prime}:\mathbb{R}^{l}\rightarrow\mathbb{R}^{k_{1}}$ such that $h(s_{1},\ldots,s_{m})=h^{\prime}(s_{t_{1}},\ldots,s_{t_{l}})$ holds for all $(s_{1},\ldots,s_{m})$. This is equivalent to $g$ depending only on $l$ order statistics of $s_{1},\ldots,s_{m}$.
We first look into the computation of $q_{L}$ and $q_{U}$ in (6). Here we assume that the rank-ordering function $\tilde{h}$ is chosen “reasonably”, so that it also depends only on the $t_{1},\ldots,t_{l}$-th components of the input. For instance, a natural choice would be
$\tilde{h}(r_{1},\ldots,r_{m})=\tilde{h}^{\prime}(r_{t_{1}},\ldots,r_{t_{l}}),\text{ where }\tilde{h}^{\prime}=h^{\prime}\big{|}_{H^{\prime}}.$
Here,
$H^{\prime}=\left\{(r_{1}^{\prime},r_{2}^{\prime},\ldots,r_{l}^{\prime}):1\leqslant r_{1}^{\prime}\leqslant\ldots\leqslant r_{l}^{\prime}\leqslant n+1\right\}.$
The first step is to compute the sizes of the level sets of the function $(r_{1},\ldots,r_{m})\mapsto(r_{t_{1}},\ldots,r_{t_{l}})$, which equal $L$ from (14). Then we compute
$L_{\tilde{h}}(\tau)=\sum_{\begin{subarray}{c}(\rho_{1},\ldots,\rho_{l}):\\ \tilde{h}^{\prime}(\rho_{1},\ldots,\rho_{l})=\tau\end{subarray}}L(\rho_{1}-1,\ldots,\rho_{l}-1)\,\text{ and }U_{\tilde{h}}(\tau)=\sum_{\begin{subarray}{c}(\rho_{1},\ldots,\rho_{l}):\\ \tilde{h}^{\prime}(\rho_{1},\ldots,\rho_{l})=\tau\end{subarray}}L(\rho_{1},\ldots,\rho_{l})$
for each $\tau\in\textrm{Im}(\tilde{h}^{\prime})$. Then, $q_{L}$ and $q_{U}$ are given by
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{\tau\in\textrm{Im}(\tilde{h}^{\prime})}\frac{L_{\tilde{h}}(\tau)}{\binom{n+m}{m}}\delta_{\tau}\right)\text{ and }q_{U}=Q_{1-\alpha}\left(\sum_{\tau\in\textrm{Im}(\tilde{h}^{\prime})}\frac{U_{\tilde{h}}(\tau)}{\binom{n+m}{m}}\delta_{\tau}\right).$
The formula for $B_{L}$ and $B_{U}$ in can be written as
$\begin{split}B_{L}&=\min\left\{h^{\prime}(S_{(r_{1}^{\prime}-1)},\ldots,S_{(r_{l}^{\prime}-1)}):(r_{1}^{\prime},\ldots,r_{l}^{\prime})\in H^{\prime},\tilde{h}^{\prime}(r_{1}^{\prime},\ldots,r_{l}^{\prime})\geqslant q_{L}\right\},\\ B_{U}&=\max\left\{h^{\prime}(S_{(r_{1}^{\prime})},\ldots,S_{(r_{l}^{\prime})}):(r_{1}^{\prime},\ldots,r_{l}^{\prime})\in H^{\prime},\tilde{h}^{\prime}(r_{1}^{\prime},\ldots,r_{l}^{\prime})\leqslant q_{U}\right\},\end{split}$ (32)
and this requires the computation of the function values at $|H^{\prime}|$ number of inputs, which scales as $n^{l}$. Therefore, we obtain a computationally feasible procedure for the case $h$ is sparse, i.e., $l$ is small.
Appendix D Additional simulation results
In this section, we reproduce the experimental results from Section 4.3 in the case where the true propensity score is unavailable, and instead, an estimate of the propensity score is used in the procedure. Specifically, we generate training data of size 200, fit a random forest classifier to construct an estimate $\hat{p}_{A\mid X}(\cdot)$ of the propensity score $p_{A\mid X}(\cdot)$, and then repeat the procedure with $p_{A\mid X}$ replaced by $\hat{p}_{A\mid X}$—i.e., we use the estimated propensity score in the rejection sampling step, and the following steps remain unchanged. The results for ths tasks of inference on the mean and quartiles are shown in Table 7 and Figure 9, illustrating that the prediction sets obtained with the estimated propensity score are similar to those from the true propensity score.
Target $\alpha=0.05$ $\alpha=0.075$ $\alpha=0.1$ $\alpha=0.125$ $\alpha=0.15$ $\alpha=0.175$ $\alpha=0.2$
Median 0.968 (0.0079) 0.952 (0.0096) 0.940 (0.0106) 0.914 (0.0126) 0.894 (0.0138) 0.870 (0.0151) 0.858 (0.0156)
Quartiles 0.968 (0.0079) 0.958 (0.0090) 0.934 (0.0111) 0.922 (0.0120) 0.902 (0.0133) 0.874 (0.0149) 0.844 (0.0162)
Table 7: Coverage rates of the batch PI prediction sets for counterfactual quartiles using the estimated propensity score (upper: median, lower: quartiles) at different levels, with standard errors.
Figure 9: Coverage rates of the batch PI prediction sets for the median and quartiles of counterfactual variables using the estimated propensity score at different levels. The dotted line corresponds to the $y=x$ line.
Next, we present the results for inference on the mean using the estimated propensity score (Table 8 and Figure 10). The results illustrate that the prediction sets obtained from the estimate still achieve the coverage guarantee, although they are a bit more conservative.
Test size $\alpha=0.05$ $\alpha=0.075$ $\alpha=0.1$ $\alpha=0.125$ $\alpha=0.15$ $\alpha=0.175$ $\alpha=0.2$
$m=5$ 0.984 (0.0056) 0.970 (0.0076) 0.956 (0.0092) 0.952 (0.0096) 0.942 (0.0105) 0.932 (0.0113) 0.926 (0.0117)
$m=10$ 0.998 (0.0020) 0.992 (0.0040) 0.986 (0.0053) 0.980 (0.0063) 0.968 (0.0079) 0.962 (0.0086) 0.950 (0.0098)
Table 8: Coverage rates of the prediction sets for the mean of counterfactual variables using the estimated propensity score for test sizes of five and ten, at different levels, along with standard errors.
Figure 10: Coverage rates of the prediction set for the mean of counterfactual variables using the estimated propensity score for test sizes five and ten, at different levels. The dotted line corresponds to $y=x$ line.
Appendix E Additional proofs
E.1 Proof of Theorem 1
We first consider the case where the scores $S_{1},\ldots S_{n},S_{n+1},\ldots,S_{n+m}$ are all distinct with probability one. By Condition 1, there exist functions $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$ and $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ such that $g(\{z_{1},\ldots,z_{m}\})=h(s(z)_{\uparrow})$ holds for any $z=(z_{1},z_{2},\ldots,z_{m})$. Recall that $S_{i}=s(X_{i},Y_{i})$ for $i\in[n+m]$ and $S_{(1)},S_{(2)},\ldots,S_{(n)}$ are the order statistics of the observed scores $S_{1},S_{2},\ldots,S_{n}$.
For $j=1,2,\ldots,m$, define
$R_{n+j}=\min\{r\in\{1,2,\ldots,n\}:S_{(r)}\geqslant S_{n+j}\},$ (33)
i.e., $R_{n+j}$ is the rank such that $S_{(R_{n+j})}$ is the smallest observed score that is larger than or equal to $S_{n+j}$. We define $R_{n+j}=n+1$ if $S_{(n)}<S_{n+j}$. Write $R^{\textnormal{test}}=(R_{n+1},R_{n+2},\ldots,R_{n+m})$. We also define $T_{i}$ as the rank (in increasing order) of $S_{i}$ among the set of all scores $\{S_{1},\ldots,S_{n},S_{n+1},\ldots,S_{n+m}\}$, for $i\in[n+m]$.
First we will show that $R_{\uparrow}^{\text{test}}\sim\textnormal{Unif}(H)$. Define the set $C_{n+m}=\left\{r_{1:m}:1\leqslant r_{1}<r_{2}<\ldots<r_{m}\leqslant n+m\right\}$, and let $T^{\text{test}}=(T_{n+1},T_{n+2},\ldots,T_{n+m})$ be the vector of ranks of the test scores. It is clear from the exchangeability of $S_{1},\ldots,S_{n+m}$ that $T_{\uparrow}^{\text{test}}$ follows a uniform distribution over $C_{n+m}$—i.e., all the rank combinations appear with the same probability. Next, we construct a map $M$ from $C_{n+m}$ to $H$ such that for all $r_{1:m}\in C_{n+m}$,
$M(r_{1:m})=(r_{1},r_{2}-1,\ldots,r_{k}-k+1,\ldots,r_{m}-m+1).$
This is a well defined function, since for any $11\leqslant k\leqslant m-$ 1, it holds that $r_{k+1}-(k+1)+1\geqslant r_{k}+1-(k+1)+1=r_{k}-k+1$. Observe that $M$ is a bijection, since it has an inverse function defined for all $r_{1:m}\in H$ by
$M^{-1}(r_{1:m})=(r_{1},r_{2}+1,\ldots,r_{k}+k-1,\ldots,r_{m}+m-1).$
Therefore, $M(T_{\uparrow}^{\text{test}})$ follows a uniform distribution over $H$.
The next step is to observe that $M(T_{\uparrow}^{\text{test}})=R_{\uparrow}^{\text{test}}$. To see this, assume $T_{n+1}<T_{n+2}<\ldots<T_{n+m}$, without loss of generality, and fix any $j\in[m]$. By the definition of $R_{n+j}$, we have
$R_{n+j}$ $=\sum_{i=1}^{n}{\mathbbm{1}}\left\{{S_{i}<S_{n+j}}\right\}+1=\sum_{i=1}^{n+m}{\mathbbm{1}}\left\{{S_{i}<S_{n+j}}\right\}-\sum_{i=n+1}^{n+m}{\mathbbm{1}}\left\{{S_{i}<S_{n+j}}\right\}+1$
$=(T_{n+j}-1)-(j-1)+1=T_{n+j}-j+1.$
Putting everything together, we have shown that $R_{\uparrow}^{\text{test}}\sim\textnormal{Unif}(H)$. This implies that, for any fixed subset $I$ of $H$ with $|I|\geqslant(1-\gamma)|H|$, it holds that $\mathbb{P}\left\{{R_{\uparrow}^{\text{test}}\in I}\right\}\geqslant 1-\gamma$. Let $S_{(n+1)},\ldots,S_{(n+m)}$ represent the order statistics of $S_{n+1},\ldots,S_{n+m}$, and $R_{(n+1)},\ldots,R_{(n+m)}$ denote the order statistics of $R_{n+1},\ldots,R_{n+m}$ (so that $R_{\uparrow}^{\text{test}}=(R_{(n+1)},\ldots,R_{(n+m)})$). Now, $S_{n+j}\leqslant S_{(R_{n+j})}$ holds for each $j\in[m]$ by the definition of $R_{n+j}$, and this implies that $S_{(n+j)}\leqslant S_{(R_{(n+j)})},j\in[m]$. Therefore, we have
$\mathbb{P}\left\{{h(S_{(n+1)},\ldots,S_{(n+m)})\leqslant\max_{r_{1:m}\in I}h(S_{(r_{1})},\ldots,S_{(r_{m})})}\right\}$
$\geqslant\mathbb{P}\left\{{h(S_{(R_{(n+1)})},\ldots,S_{(R_{(n+m)})})\leqslant\max_{r_{1:m}\in I}h(S_{(r_{1})},\ldots,S_{(r_{m})})}\right\}\geqslant\mathbb{P}\left\{{(R_{(n+1)},\ldots,R_{(n+m)})\in I}\right\}\geqslant 1-\gamma,$
where the first inequality applies the monotonicity assumption (3) of $h$ and the definition of $R_{n+1},\ldots,R_{n+m}$, and the second inequality uses the inclusion $\{f(x)\leqslant\max_{y\in A}f(y)\}\supset\{x\in A\}$, valid for any function $f$ defined on a finite set $B$, for any $A\subset B$ and any $x\in B$. Further, $B_{U}=\max_{r_{1:m}\in I}h(S_{(r_{1})},\ldots,S_{(r_{m})})$ where $I:=\left\{r_{1:m}\in H,\tilde{h}(r_{1:m})\leqslant q_{U}\right\}$. Since $|I|\geqslant(1-\gamma)|H|$ by the definition of $q_{U}$, we have $\mathbb{P}\left\{{h(S_{(n+1)},\ldots,S_{(n+m)})\leqslant B_{U}}\right\}\geqslant 1-\gamma.$
For the lower bound, we first observe that $S_{(R_{n+j}-1)}<S_{n+j}$ for each $j\in[m]$, by the definition of $R_{n+j}$. Then $S_{(R_{(n+j)}-1)}<S_{(n+j)}$ also holds, and thus
$h(S_{(n+1)},\ldots,S_{(n+m)})\geqslant h(S_{(R_{(n+1)}-1)},\ldots,S_{(R_{(n+m)}-1)})$
holds deterministically. Thus, following an argument similar to that for the upper bound, we can prove that $\mathbb{P}\left\{{h(S_{(n+1)},\ldots,S_{(n+m)})\geqslant B_{L}}\right\}\geqslant 1-\beta$ also holds, and this proves the desired inequality.
Now consider the case where the scores can have ties. In such a case, we define $\tilde{T}_{i}$ as the rank of $S_{i}$ among $\{S_{1},S_{2},\ldots,S_{n+m}\}$, where we break the ties uniformly randomly. For example, if $S_{2}<S_{1}=S_{3}<S_{4}$, then we have $T_{2}=1,T_{4}=4$ deterministically, and $(T_{2},T_{3})=(2,3)$ and $(T_{2},T_{3})=(3,2)$ each with probability $1/2$. Let $\tilde{T}_{(1)}^{\text{cal}}<\ldots<\tilde{T}_{(n)}^{\text{cal}}$ be the order statistics of $\{\tilde{T}_{i}:i\in[n]\}$. Then we let
$\tilde{R}_{n+j}=\min\{r\in[n]:\tilde{T}_{(r)}^{\text{cal}}\geqslant T_{n+j}\}.$
By the same argument as before, we have that $\tilde{R}_{\uparrow}^{\text{test}}=(\tilde{R}_{(n+1)},\ldots,\tilde{R}_{(n+m)})\sim\textnormal{Unif}(H)$. Also note that $\tilde{R}_{n+j}\geqslant R_{n+j}$ holds for all $j\in[m]$, since $\tilde{T}_{(r)}^{\text{cal}}\geqslant T_{n+j}$ implies $S_{(r)}\geqslant S_{n+j}$ (i.e., $\tilde{T}_{(r)}^{\text{cal}}\geqslant T_{n+j}$ cannot happen if $S_{(r)}<S_{n+j}$). Therefore, we have $\tilde{R}_{\uparrow}^{\text{test}}\succeq R_{\uparrow}^{\text{test}}$, and thus it follows that
$\mathbb{P}\left\{{h(S_{(n+1)},\ldots,S_{(n+m)})\leqslant\max_{r_{1:m}\in I}h(S_{(r_{1})},\ldots,S_{(r_{m})})}\right\}$
$\geqslant\mathbb{P}\left\{{h(S_{(R_{(n+1)})},\ldots,S_{(R_{(n+m)})})\leqslant\max_{r_{1:m}\in I}h(S_{(r_{1})},\ldots,S_{(r_{m})})}\right\}$
$\geqslant\mathbb{P}\left\{{h(S_{(\tilde{R}_{(n+1)})},\ldots,S_{(\tilde{R}_{(n+m)})})\leqslant\max_{r_{1:m}\in I}h(S_{(r_{1})},\ldots,S_{(r_{m})})}\right\}\geqslant\mathbb{P}\left\{{(\tilde{R}_{(n+1)},\ldots,\tilde{R}_{(n+m)})\in I}\right\}\geqslant 1-\gamma,$
proving the claim.
E.2 Proof of Proposition 1
Given non-negative integers $\delta_{1},\ldots,\delta_{m}$, define $r_{i}=\sum_{j\in[i]}\delta_{j}$ for all $i\in[m]$. Further, for any $n\geqslant r_{m}$, recalling $\delta_{1:m}=(\delta_{1},\ldots,\delta_{m})$ define $g$ via $g(\delta_{1:m})=h(S_{(r_{1})},\ldots,S_{(r_{m})})$. Clearly, the constraint $r_{1:m}\in H$ holds. Choosing $\tilde{h}\equiv 0$, $B_{L}$ from (7) becomes
$\min\left\{g(\delta_{1:m}):\delta_{i}\in\{0,\ldots,n\},i\in[m],\,\sum_{j\in[m]}\delta_{j}\leqslant n\right\}.$
By taking $g$ to take sufficiently large polynomial-sized values when any $\delta_{i}\geqslant 2$, $i\in[m]$, we can constrain $\delta_{i}\in\{0,1\},i\in[m]$. Further, we can take $n=m$. Since $g$ can be arbitrary, we now claim that the above problem includes the vertex cover problem [see e.g., Garey and Johnson, 1979] as a special case.
Indeed, given a graph $G=(V,E)$ and $\lambda\in\mathbb{R}$, we can take $g$ to be $g(\delta_{1:m})=\sum_{u\in V}\delta_{u}+\lambda\sum_{(u,v)\in E}(1-\delta_{u}-\delta_{v})_{+}$ for $\delta_{1:m}\in\{0,1\}^{m}$, where $(\cdot)_{+}$ is the positive part. Next, we claim that for $\lambda\leqslant|V|+1$, any minimizer $(\delta_{1:m})$ of $g$ must satisfy $\delta_{u}+\delta_{v}\geqslant 1$ for all $(u,v)\in E$. Indeed, otherwise $\lambda\sum_{(u,v)\in E}(1-\delta_{u}-\delta_{v})_{+}\geqslant\lambda$; whereas setting $\tilde{\delta}_{u}=1$ for all $u\in V$ leads to a value of $g(\tilde{\delta}_{1},\ldots,\tilde{\delta}_{m})=|V|<\lambda$; which is a contradiction with $(\delta_{1},\ldots,\delta_{m})$ being a minimizer.
Now, a minimizer of $\sum_{u\in V}\delta_{u}$ with $\delta_{u}\in\{0,1\}$ for all $u\in V$ and $\delta_{u}+\delta_{v}=1$ for all $(u,v)\in E$ exists and corresponds to a vertex cover; and all such minimizers are vertex covers. This shows that for this $\lambda$, the minimizers of $g$ are precisely the vertex covers. We conclude that our problem includes the vertex cover problem as a special case, and hence is NP-hard.
E.3 Proof of Theorem 2
Let us define $R_{n+1},R_{n+2},\cdots,R_{n+m}$ as in (33). Then, it holds that
$\mathbb{P}\left\{{S_{(w_{1}-1)}\leqslant S_{(t_{1})}^{\textnormal{test}}\leqslant S_{(q_{1})},\cdots,S_{(w_{l}-1)}\leqslant S_{(t_{l})}^{\textnormal{test}}\leqslant S_{(q_{l})}}\right\}\\ \geqslant\mathbb{P}\left\{{S_{(w_{1})}\leqslant S_{(R_{n+t_{1}})}\leqslant S_{(q_{1})},\cdots,S_{w_{l}}\leqslant S_{(R_{n+t_{l}})}\leqslant S_{(q_{l})}}\right\}\\ \geqslant\mathbb{P}\left\{{w_{1}\leqslant R_{n+t_{1}}\leqslant q_{1},\cdots,w_{l}\leqslant R_{n+t_{l}}\leqslant q_{l}}\right\}\geqslant 1-\alpha,$
where the last inequality holds by the condition $F_{n,m}(w_{1},\cdots,w_{l};q_{1},\cdots,q_{l})\geqslant(1-\alpha)\cdot|H|$ and the fact that $R_{\uparrow}^{\text{test}}\sim\textnormal{Unif}(H)$ holds by the result in the proof of Theorem 1.
E.4 Proof of Corollary 2
The proof follows directly from the definition of $B$ in (16) and Theorem 2.
E.5 Proof of Proposition 2
Fix any $z_{1},\ldots,z_{n},z_{n+1},\ldots,z_{n+m}$, where each $z_{i}=(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$, and let $\mathcal{E}_{z}$ denote the event that $\{Z_{1},\ldots,Z_{n},Z_{n+1},\ldots,Z_{n+m}\}=\{z_{1},\ldots,z_{n},z_{n+1},\ldots,z_{n+m}\}$, indicating that the data points are equal to these specified values as a (multi-)set. For simplicity, let us also write $\mathcal{E}_{A}$ to denote the event $A_{1}=\ldots=A_{n}=1,A_{n+1}=\ldots=A_{n+m}=0$.
Let $\mathcal{S}_{n+m}$ denote the set of all permutations of $[n+m]$. For $I=\{i_{1},\ldots,i_{m}\}$ with $1\leqslant i_{1}<\ldots<i_{m}\leqslant n$, we compute
$\mathbb{P}\left\{{\{Z_{n+1},\ldots,Z_{n+m}\}=\{z_{i_{1}},\ldots,z_{i_{m}}\}}\ \middle|\ {\mathcal{E}_{z},\mathcal{E}_{A}}\right\}$
=$=\frac{\mathbb{P}\left\{{\mathcal{E}_{A}}\ \middle|\ {\{Z_{n+1},\ldots,Z_{n+m}\}=\{z_{i_{1}},\ldots,z_{i_{m}}\},\mathcal{E}_{z}}\right\}\cdot\mathbb{P}\left\{{\{Z_{n+1},\ldots,Z_{n+m}\}=\{z_{i_{1}},\ldots,z_{i_{m}}\}}\ \middle|\ {\mathcal{E}_{z}}\right\}}{\mathbb{P}\left\{{\mathcal{E}_{A}}\ \middle|\ {\mathcal{E}_{z}}\right\}}$
$=\frac{\prod_{k=1}^{m}(1-p_{A\mid X})(x_{i_{k}})\cdot\prod_{i\notin\{i_{1},\ldots,i_{m}\}}p_{A\mid X}(x_{i})\cdot\frac{n!m!}{(n+m)!}}{\sum_{\sigma\in S_{n+m}}\mathbb{P}\left\{{\mathcal{E}_{A},Z_{n+1}=z_{\sigma(1)},\ldots,Z_{n+m}=z_{\sigma(n+m)}}\ \middle|\ {\mathcal{E}_{z}}\right\}}$
$=\frac{\prod_{k=1}^{m}(1-p_{A\mid X}(x_{i_{k}}))\cdot\prod_{i\notin\{i_{1},\ldots,i_{m}\}}p_{A\mid X}(x_{i})\cdot\frac{n!m!}{(n+m)!}}{\sum_{\sigma\in\mathcal{S}_{n+m}}\frac{1}{(n+m)!}\prod_{i=1}^{n}p_{A\mid X}(x_{\sigma(i)})\cdot\prod_{i=n+1}^{n+m}(1-p_{A\mid X}(x_{\sigma(i)}))}.$
By dividing both the numerator and the denominator by $\prod_{i=1}^{n+m}p_{A\mid X}(x_{i})$, we find that this further equals
$\frac{n!m!\prod_{k=1}^{m}\frac{1-p_{A\mid X}(x_{i_{k}})}{p_{A\mid X}(x_{i_{k}})}}{\sum_{\sigma\in\mathcal{S}_{n+m}}\prod_{k=1}^{m}\frac{1-p_{A\mid X}(x_{\sigma(i)})}{p_{A\mid X}(x_{\sigma(i)})}}=\frac{n!m!\prod_{k=1}^{m}\frac{1-p_{A\mid X}(x_{i_{k}})}{p_{A\mid X}(x_{i_{k}})}}{\sum_{I\subset[n+m],|I|=m}\sum_{\sigma\in\mathcal{S}_{n+m}:\{\sigma(k):k\in[m]\}=I}\prod_{i\in I}\frac{1-p_{A\mid X}(x_{i})}{p_{A\mid X}(x_{i})}}$
$=$ $\frac{\prod_{k=1}^{m}\frac{1-p_{A\mid X}(x_{i_{k}})}{p_{A\mid X}(x_{i_{k}})}}{\sum_{I\subset[n+m],|I|=m}\prod_{i\in I}\frac{1-p_{A\mid X}(x_{i})}{p_{A\mid X}(x_{i})}}(=:p_{A\mid X}^{z}(I)).$
Therefore, we have
$g(\{Z_{n+1},\ldots,Z_{n+m}\})\mid\mathcal{E}_{z},\mathcal{E}_{A}\sim\sum_{I\subset[n+m],|I|=m}p_{A\mid X}^{z}(I)\cdot\delta_{h(S_{I}^{z})},$
where $S_{I}^{z}=(s(z_{i_{1}}),s(z_{i_{2}}),\ldots,s(z_{i_{m}}))$. It follows that
$\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\leqslant Q_{1-\gamma}\left(\sum_{I\subset[n+m],|I|=m}p_{A\mid X}^{z}(I)\cdot\delta_{h(S_{I}^{z})}\right)}\ \middle|\ {\mathcal{E}_{z},\mathcal{E}_{A}}\right\}\geqslant 1-\gamma,$
and marginalizing with respect to $\mathcal{E}_{z}$ yields $\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\leqslant Q_{1-\gamma}\left(\sum_{I\subset[n+m],|I|=m}p_{A\mid X}(I)\cdot\delta_{h(S_{I}^{Z})}\right)}\ \middle|\ {\mathcal{E}_{A}}\right\}\geqslant 1-\gamma$. By the monotonicity assumption of $h$, $h(S_{I}^{Z})\leqslant h(\overline{S}_{I})$ holds deterministically, leading to
$\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\leqslant Q_{1-\gamma}\left(\sum_{I\subset[n+m],|I|=m}p_{A\mid X}(I)\cdot\delta_{h(\overline{S}_{I})}\right)}\ \middle|\ {\mathcal{E}_{A}}\right\}\geqslant 1-\gamma.$
Similarly, we obtain $\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\geqslant Q_{\beta}^{\prime}\left(\sum_{I\subset[n+m],|I|=m}p_{A\mid X}(I)\cdot\delta_{h(\underline{S}_{I})}\right)}\ \middle|\ {\mathcal{E}_{A}}\right\}\geqslant 1-\beta$, and the desired inequality follows.
E.6 Proof of Corollary 3
It is sufficient to show that the random variables in the set $\widetilde{\mathcal{D}}_{n}\cup\{(X_{i},Y_{i}):n+1\leqslant i\leqslant n+m\}$ are i.i.d. conditional on $B_{1:n}$. Since each outcome $Y_{i}$ depends only on $X_{i}$ (i.e., independent of every other random variable conditional on $X_{i}$) and is drawn from the same distribution $P_{Y\mid X}$, it is further enough to show that $\{X_{i}:i\in[n],B_{i}=1\}\cup\{X_{i}:n+1\leqslant i\leqslant n+m\}$ are i.i.d. given $B_{1:n}$. The independence is clear under the model (19), and thus it remains to prove that the following two distributions are identical.
  1. 1.
    Conditional distribution of $X$ given $B=1$, where $X$ and $B$ are drawn by $X\sim P_{X\mid A=1},B\mid X\sim\textnormal{Bern}(p_{B\mid X}(X))$.
  2. 2.
    The distribution $P_{X\mid A=0}$.
Take any measurable set $U\subset\mathcal{X}$. We have
$\mathbb{P}_{{X\sim P_{X\mid A=1},B\mid X\sim\textnormal{Bern}(p_{B\mid X}(X))}}\left\{{X\in U}\ \middle|\ {B=1}\right\}$
$=\mathbb{P}_{{X\sim P_{X},A\mid X\sim\textnormal{Bern}(p_{A\mid X}(X)),B\mid X\sim\textnormal{Bern}(p_{B\mid X}(X))}}\left\{{X\in U}\ \middle|\ {B=1,A=1}\right\}$
=$=\frac{\mathbb{P}\left\{{A=1,B=1}\ \middle|\ {X\in U}\right\}\cdot\mathbb{P}\left\{{X\in U}\right\}}{\mathbb{P}\left\{{A=1,B=1}\right\}}=\frac{\mathbb{E}\left[{\mathbb{P}\left\{{A=1,B=1}\ \middle|\ {X}\right\}}\ \middle|\ {X\in U}\right]\cdot\mathbb{P}\left\{{X\in U}\right\}}{\mathbb{E}\left[{\mathbb{P}\left\{{A=1,B=1}\ \middle|\ {X}\right\}}\right]}$
$=\frac{\mathbb{E}\left[{p_{A\mid X}(X)\cdot\frac{c}{1-c}\cdot\frac{1-p_{A\mid X}(X)}{p_{A\mid X}(X)}}\ \middle|\ {X\in U}\right]\cdot\mathbb{P}\left\{{X\in U}\right\}}{\mathbb{E}\left[{p_{A\mid X}(X)\cdot\frac{c}{1-c}\cdot\frac{1-p_{A\mid X}(X)}{p_{A\mid X}(X)}}\right]}=\frac{\mathbb{E}\left[{1-p_{A\mid X}(X)}\ \middle|\ {X\in U}\right]\cdot\mathbb{P}\left\{{X\in U}\right\}}{\mathbb{E}\left[{1-p_{A\mid X}(X)}\right]}$
$=\frac{\mathbb{E}\left[{\mathbb{P}\left\{{A=0}\ \middle|\ {X}\right\}}\ \middle|\ {X\in U}\right]\cdot\mathbb{P}\left\{{X\in U}\right\}}{\mathbb{E}\left[{\mathbb{P}\left\{{A=0}\ \middle|\ {X}\right\}}\right]}=\frac{\mathbb{P}\left\{{A=0}\ \middle|\ {X\in U}\right\}\cdot\mathbb{P}\left\{{X\in U}\right\}}{\mathbb{P}\left\{{A=0}\right\}}=\mathbb{P}\left\{{X\in U}\ \middle|\ {A=0}\right\}$
$=\mathbb{P}_{{X\sim P_{X\mid A=0}}}\left\{{X\in U}\right\}.$
This shows that the above two distributions are identical, and thus the claim is proved.
E.7 Proof of Corollary 4
By Theorem 1 and the observations in Section 2.2.1 for inference on the quantile, we have $\mathbb{P}\left\{{S_{(m-\eta)}^{\textnormal{test}}\leqslant\hat{T}}\right\}\geqslant 1-\alpha$. Now, the event $\{S_{n+j}=\hat{\mu}(X_{n+j}){\mathbbm{1}}\left\{{Y_{n+j}\leqslant c}\right\}>\hat{T}\}$ is equivalent to the event $\{\hat{\mu}(X_{n+j})>\hat{T}\text{ and }Y_{n+j}\leqslant c\}$, since $\hat{T}\geqslant 0$ holds almost surely. Therefore,
$\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})>\hat{T},Y_{n+j}\leqslant c}\right\}\leqslant\eta}\right\}=\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{S_{n+j}>\hat{T}}\right\}\leqslant\eta}\right\}=\mathbb{P}\left\{{S_{(m-\eta)}^{\textnormal{test}}\leqslant\hat{T}}\right\}\geqslant 1-\alpha,$
as desired.